### v2

Abstract
Choice overload – by which larger choice sets are detrimental to a chooser’s wellbeing – is potentially of great importance to the design of economic policy. Yet the current evidence on its prevalence is inconclusive. We argue that existing tests are likely to be underpowered and hence that choice overload may occur more often than the literature suggests. We propose more powerful tests based on richer data and characterization theorems for the Random Utility Model. These new approaches come with significant econometric challenges, which we show how to address. We apply our tests to new experimental data and find strong evidence of choice overload that would likely be missed using current approaches.
1 Introduction
The standard model of utility maximization tells us that increasing the set of available options can make the consumer no worse off, and may well improve their welfare. Since the pioneering study of Iyengar and Lepper (2000), [^1] See also Reibstein et al. (1975). a body of work in psychology has called this assumption into question. The umbrella term “choice overload” covers a number of phenomena by which larger choice sets seem to make people worse off. By now there are a huge number of studies investigating various aspects of choice overload; see Scheibehenne et al. (2010) and Chernev et al. (2015) for recent reviews and meta analyses.
Some of the output measures used to identify choice overload are hard to interpret using the classic tools of economic analysis; examples include a reduction in ex-post reported satisfaction or lower confidence that the right choice was made. Others fall very much in the realm of choice theory. In this paper we focus on the observation that larger choice sets may make people more likely to choose a “default” option.
Iyengar and Lepper (2000) provide an example of this type of behavior in their famous “jam” study. On different days a table was set up in a supermarket displaying a range of jams and offering a money-off coupon if a jam was bought. On “limited” choice days, six jams were available; on “extensive” choice days, customers faced 18 additional options. Strikingly, people were less likely to buy a jam on extensive choice days (2% vs 12% on limited choice days). [^2] These are the unconditional probabilities of buying jam. Typically the numbers quoted for this study are the probabilities of buying jam conditional on stopping at the table (3% vs 30%). However these numbers are more difficult to interpret as there may be selection into who stops at the table. Treating “do not buy” as the default choice option, this violates the Independence of Irrelevant Alternatives (IIA) axiom.
If choice overload could be well established it would have a number of important implications for economics. From a positive perspective, it would clearly violate the standard model of utility maximization. Moreover, it is not well explained by some of the more obvious behavioral theories. For example, most models of reference dependence assume that, for choices that share the same reference point, the decision maker will behave like a standard utility maximizer, ruling out choice overload type effects (Dean, 2008). Models of limited attention can also struggle to replicate the phenomenon, if they cannot explain why an overloaded individual does not simply ignore some of the available options. [^3] Though see Cattaneo et al. (2020, 2021) for examples of models of consideration sets, and de Lara and Dean (2024) for a model of sequential search that can accommodate choice overload-type behavior. This points to the need for models including more exotic elements such as regret (Sarver, 2008), rational contextual inference (Kamenica, 2008; Nocke and Rey, 2021) or decision avoidance (Beattie et al., 1994; Dean, 2008; Gerasimou, 2018). From a normative perspective, a key tenet of classic economics is that welfare is weakly improved by increasing the set of options available to the consumer. This assumption forms the basis of many policy recommendations, yet choice overload would call it into question.
Unfortunately, current research into choice overload is inconclusive. Some direct replications of previous experiments have failed (Scheibehenne, 2008; Greifeneder et al., 2010). One recent meta-analysis concluded that the mean measured choice overload effect is zero (Scheibehenne et al., 2010). Another one (Chernev et al., 2015) concludes that whether or not choice overload exists may depend a lot on context.
The aim of this paper is to argue that existing studies are likely to underestimate the degree of choice overload of this type. [^4] Despite the above discussion, from now on we will use ‘choice overload’ to refer specifically to the phenomenon of choosing a default option more often in larger choice sets. Almost universally, a data set is said to exhibit choice overload only if the default is more likely to be chosen in a larger choice set than it is in some smaller choice set (see section 2). Yet intuitively, under the hypothesis of utility maximization one would expect the default to be chosen much less in larger choice sets, because the decision maker (DM) should be more likely to find something that they prefer to the default. This lack of power may potentially explain the inconsistent results from existing tests.
Two examples may further clarify this point.
Example 1.1.
[…]
Example 1.2.
The following example is even more stylized but avoids distributional assumptions. Let the choice universe be $X=\{a,b,c,d\}$, where $d$ stands for default. Suppose $a,b,c$ are individually chosen with probability $20\%$ from choice sets $\{a,d\},\{b,d\},\{c,d\}$. Suppose also that the probability of active choice is $40\%$ for any of $\{a,b,d\},\{a,c,d\},\{b,c,d\}$. Assuming that indifference is not allowed, these probabilities can be reconciled with a Random Utility Model (RUM, precisely defined later); however, they imply that the preferences $a\succ d$, $b\succ d$, and $c\succ d$ occur in nonoverlapping subsets of the population. Thus, the active choice probability from $X$ must be at least $60\%$. If it is lower, we observe choice overload in the sense that observations cannot be explained in terms of a fully rational population, but they can be explained in terms of a population that is fully rational except that some individuals switch to default choice when encountering the grand choice set. This is true even if the active choice probability from $X$ strictly exceeds the active choice probability from any smaller set.
In this paper, we describe new tests for choice overload. We consider a data set based on a finite set of alternatives $X$ that includes a default $d$. Observations of choice behavior are made from the grand set $X$, plus some subsets that always contain $d$. In each case we assume that the probability that $d$ is chosen is observed. We make use of two definitions of choice overload. The first is model free: we say that a data set exhibits choice overload if it violates monotonicity – i.e., the probability of choosing the default option increases as more alternatives are added to the choice set. The second is based on the null hypothesis of the random utility model (RUM) – i.e. there is a set of utility functions, a (choice set independent) probability distribution over that set, and the probability of choosing any option in any choice set is the probability of the set of utility functions that make that option maximal. As example 1.2 illustrates, consistency with RUM implies monotonicity, but not vise versa, so the latter definition provides a more sensitive test for choice overload. [^5] In general, it is well known that stochastic monotonicity is necessary but not sufficient for random utility. Example 1.2 clarifies that coarsening the domain of observations to active versus passive choice does not fundamentally change this. The advantage of the former is that it does not rely on the underlying assumption of random utility - monotonicity is also implied by a number of other models of stochastic choice, such as Manzini and Mariotti’s (2014) Random Consideration Set model.
If the data set consists only of choice probabilities from $X$ and a single subset $A$ (as is the case in Iyengar and Lepper (2000) and most other studies), then both definitions collapse to the conventional test: Assuming that only the probability of default choice is observed, the data set is consistent with monotonicity and RUM if and only if the probability of choosing $d$ from $X$ is weakly lower than from $A$. We therefore propose as a first modification of the standard approach that data should be collected from multiple subsets of $X$.
In this richer data, monotonicity implies that the choice of the default from $X$ is lower than the lowest probability of choosing $d$ across all observed choice sets $A$. We call this the Min bound. While easy to define, testing the Min bound is subtle because it is a test of multiple hypotheses. Even if monotonicity holds, if a finite number of observations are collected from many subsets there is a high probability that one will exhibit lower default choice than is observed from $X$. Essentially, the problem is the same as that of the “winners curse” – because the probability of default choice is observed with noise, the minimum of many observations is likely to be below the true minimum. To deal with this, we propose tests that take inspiration from closely related problems in the literature on moment inequalities.
Because monotonicity is necessary but not sufficient for RUM, our model-based definition of choice overload allows us to identify tighter bounds. We adapt the approach of McFadden and Richter (1991), who show that random choice data can be rationalized by the RUM if, and only if, a particular linear expression has a solution. We refer to the highest default choice probability in the large set for which such a solution exists as the RUM bound. The fact that it is choice frequencies, not underlying probabilities, that are observed presents a significant econometric challenge. We adapt the nonparametric test of RUM proposed by Kitamura and Stoye (2018), potentially including computational innovations by Smeulders et al. (2020) needed for applications with larger sets $X$.
We apply our tests to a novel experimental data set based on choices from subsets of size $2$ and $3$ of $12$ different choice objects plus a default. The objects are verbally described sums of numbers as used in Caplin et al. (2011); see Figures 1 and 2 for a preview. In any given choice set, the subject could choose to stick with a default which provided \$3.50 for sure, or could switch to other available choices. Subjects were asked to chose from $10$ such sets, one of which was the default plus the entire set of $12$ other options –henceforth called the grand set– while the other $9$ were uniformly randomly selected from all possible sets with $1$ or $2$ options plus default.
A total of $2000$ subjects were recruited through Amazon’s Mechanical Turk platform. After removing subjects who failed a comprehension quiz, we were left with $1832$ subjects and $18320$ choices.
Traditional approaches would be unlikely to find evidence of choice overload in our data set: the average default choice in small sets is higher than in the grand set, and only 15 of the 78 small sets have a default choice which is significantly lower than that of the grand set. Hence if, as in most work in the literature, data was collected from only a single or few small sets, one would be unlikely to detect overload. In contrast, even the most ‘brute force’ of our proposed tests –a potentially very conservative but finite-sample valid implementation of the Min bound– do detect it.
Our test based on RUM provides an unexpected additional insight: While RUM is indeed rejected on the whole data set, it is also rejected if we remove data from choices in the grand set. This is because subjects are more likely to choose the default option in choice sets of size $3$ than their data from size $2$ sets and RUM would predict. Thus, our results indicate that choice overload type effects can start in choice sets which are much smaller than have previously been demonstrated.
We hope that our work will have three consequences. First, by providing a higher powered test of choice overload, it should clear up the question of whether this is indeed a real phenomenon. Second, given that (we suspect) it will show choice overload to be more prevalent than previously thought, we hope it will spur further theoretical and policy work designed to understand its causes and mitigate its effects. Finally, by providing a better tool for measuring when choice overload does occur, we hope it will facilitate the above work by providing a better empirical basis on which to theorize.
2 Literature Review
Many studies have examined the phenomenon of choice overload from different perspectives, including different outcome measures and potential moderators. Two recent meta-analyses (Chernev et al., 2015; Scheibehenne et al., 2010) report results from 99 experiments in 53 papers and 63 experiments in 50 papers, respectively. Given these two reviews, our aim here is not to provide comprehensive coverage of the existing research. Instead, we make two points.
First, to the best of our knowledge, no other paper has made use of the methodology we propose. Using the aforecited meta-analyses and Google Scholar, we identified 32 studies from 19 papers that use default choice as a measure of choice overload. Of these, 20 ask subjects to make choices from a single subset of the grand choice set. These studies can do no better than to compare the default choice probability in the small and large choice set. The remaining 12 studies ask subjects to make choices from multiple subsets of the large choice set, and so in principle could have implemented either the min bound or tests based on the RUM. However, all of them instead compare the average default choice across all small choice sets to that in the larger set, a measure which is necessary but sufficient for the aforementioned Min bound, which in turn is necessary but not sufficient for consistency with RUM. Applying our approach to the data from these previous studies (e.g., Chernev (2005), Gingras (2003), or White and Hoffrage (2009)) is an interesting avenue for future research.
While not explicitly designed to test for choice overload, the experiment of Aguiar et al. (2023) does contain the type of data needed to perform our test. Choices were observed from all subsets of a grand set of six lotteries, with a default that is always available. The authors report finding no evidence of choice overload, and our tests confirm this result. [^6] While we performed the test on all their treatments, their ‘low’ complexity treatment is the best fit to apply the test. This is because, in the ‘medium’ and ‘high’ treatments, lottery prizes were expressed as sums, and different subjects saw different sums, arguably making them different choice alternatives. This may be due to the fact that the default alternative was chosen to be ‘obviously’ worse than the other available alternatives.
The second point is that the veracity and scope of choice overload is far from established. Some direct replications have failed (Scheibehenne, 2008). One meta-analysis (Scheibehenne et al., 2010) finds a mean effect of set size on measures of choice overload to be zero, but noted a high degree of variance. A more recent analysis (Chernev et al., 2015) identifies four variables which can increase the incidence of choice overload: decision difficulty (for example due to time constraints), choice set complexity (for example due to hard to compare alternatives), preference uncertainty (for example because the decision maker is unsure how to aggregate their preferences across many dimensions), and decision goal (for example because the decision maker is not really committed to making a purchase). Chernev et al. (2015) argue that, taking these mediators into account, there is evidence of a robust choice overload effect.
We believe that the measures we introduce can contribute to this ongoing debate by providing more highly powered evidence on when choice overload is occurring. Low powered tests may in part explain some of the contradictory results in the existing literature – for example, some studies may by chance use a small choice set that leads to choice overload, while others do not. By making it easier to spot cases of overload, we anticipate movement towards a greater acceptance that choice overload does indeed occur.
3 Theory
We now present the theoretical underpinnings of our test for choice overload. We do so in three stages. First, we describe our tests under the assumption that we can perfectly observe the probabilities with which each alternative is chosen in each choice set. Second, we discuss the econometric techniques needed to handle the fact that our experimental data are sample frequencies, not population probabilities. Finally, we describe the steps necessary to implement these econometric tests, with additional considerations that proved unnecessary for our specific experiment relegated to Appendix B.
3.1 The Population-Level Testing Problem
Let $X$ be a finite set of alternatives and $d$ a default alternative contained in $X$. Let $\mathcal{D}\subset 2^{X}/\emptyset$ be a collection of choice sets, all of which contain the default alternative, and which includes the grand set $X$.
Our data set consists of observations of choice behavior in each set in $\mathcal{D}$ [^7] We follow established terminology in revealed preference theory by calling this a “data set,” but note that statistically speaking, these are population probabilities; inference for finite samples is considered in the next subsection. Specifically, we assume that we observe a function $p_{d}:\mathcal{D}\rightarrow[0,1]$, with the interpretation that $p_{d}(A)$ is the probability of choosing the default $d$ from choice set $A$.
Two notes on our data are in order. First, we assume that we observe the population probability with which the default is chosen in each choice set, but not that we can track individuals across choice problems; i.e., we observe a repeated cross section rather than a panel. This makes our approach applicable to many between-subject data sets that have this property. [^8] We note that the data we collect in our experiment is somewhat richer than this, as it contains multiple observations from the same individual. We maintain this assumption to increase the generalizability of our approach. Second, we assume that we observe only the probability with which the default was chosen in each choice set, not the probability with which specific non-default options are chosen. This is consistent with our desire to to focus on choice overload effects; the limited data only allows us to detect violations of utility maximization that occur to due too much or too little default choice from a set.
In order to design a test for choice overload, we first need to precisely define the term. Here we take two approaches. The first is to use an explicitly model-free definition, in the same way that risk aversion can be defined without reference to expected utility. An obvious candidate for such a definition is through violations of choice monotonicity, which states that the probability of choosing $d$ cannot increase as more options are added to the choice set.
Definition 3.1.
A data set $\{\mathcal{D},p\}$ satisfies monotonicity if, for any $A,B\in\mathcal{D}$ such that $A\subset B$,
$p_{d}(A)\geq p_{d}(B)$
Specifically, one could declare choice overload to have occurred if there is a higher probability of default choice in the large choice set than in the smaller one. The canonical choice overload experiment, in which $\mathcal{D}=\{A,X\}$, applies this test.
More generally, one can define
$p_{d}^{min}(X)=\min_{A\in\mathcal{D}\setminus X}p_{d}(A)$
as the smallest observed probability of choosing $d$ in any set other than $X$. If $p_{d}(X)>p_{d}^{min}(X)$ then monotonicity is violated and the data set exhibits choice overload. We refer to $p_{d}^{min}(X)$ as the Min bound. Data that violates the Min bound is inconsistent with a number of models – most obviously RUM, but also the stochastic consideration set model of Manzini and Mariotti (2014) [^9] Other, more general models of stochastic consideration do allow for choice overload type effects - see for example Cattaneo et al. (2020, 2021) and models of reference dependent preferences such as Tversky and Kahneman (1991).
A second approach is to define choice overload as a violation of a specific model. Here, the most obvious candidate would be RUM, and we will work with it, although the basic idea would generalize to any model that we know how to test. Thus, call a data set stochastically rationalizable if it could have been generated by a RUM. Then we can think of it as revealing choice overload if it would be stochastically rationalizable except that the probability of default choice in larger choice sets is too large.
To illustrate this idea, consider the following definitions. Note that, while $\mathcal{D}$ in our application contains $X$, these definitions refer to a generic $\mathcal{D}$ that may or may not do so.
Definition 3.2.
A data set $\{\mathcal{D},p\}$ is consistent with RUM if there exist a finite collection $\mathcal{U}$ of one-to-one utility functions on $X$ and a probability distribution $\rho\in\Delta(\mathcal{U})$ such that, for every $A\in\mathcal{D}$,
$p_{d}(A)=\sum_{u\in\mathcal{U}\mid d=\arg\max u(A)}\rho(u)$
For any set $A\subseteq X$, we can then define a maximal bound on the choice of default using the default choice probabilities from the subsets of $A$ and consistency with RUM. The basic idea here goes back to Varian (1982, 1983): A counterfactual choice behavior is in the predictive bounds if, and only if, that choice behavior and previously observed ones are jointly rationalizable. Formally, we have:
Definition 3.3.
For a dataset $\{\mathcal{D},p\}$ and choice problem $A\subseteq X$, define
$p_{d}^{RUM}(A)=\sup x\in[0,1]$
such that the data set
$\tilde{p}_{d}(\tilde{A})=\left\{\begin{array}[]{c}x\text{ if }\tilde{A}=A\\ p_{d}(\tilde{A})\text{ otherwise}\end{array}\right.$
defined on the choice sets $\mathcal{D}_{A}\equiv\{\tilde{A}\in\mathcal{D}:\tilde{A}\subset A\}$ is consistent with RUM.
For most of this paper, we say that a data set exhibits choice overload if $p_{d}(X)>p^{RUM}_{d}(X)$. However, the definition of $p_{d}^{RUM}(\cdot)$ allows for choice overload to “kick in” for smaller choice sets and we will consider that later.
Conceptually, the Min bound and the RUM bound deal with different reasons why observations from a single small choice set might not effectively identify choice overload. The first is heterogeneity in the quality of available items. Consider a grand choice set that consists of a number of not very appealing jam flavors (gooseberry, bark, salmon) and one extremely appealing flavor (strawberry). Absent any choice overload, we would expect to see high levels of default choice in small sets that do not include the strawberry jam, and low levels of default choice in both small sets that included the strawberry jam, and the grand set. Thus, if a researcher randomly selected for analysis a small choice set without the strawberry jam, it would make it very hard to spot choice overload. Collecting data on all small choice sets and applying the Min bound would help to solve this problem, because the researcher would observe choice from a small set in which strawberry jam was included.
A second issue is preference heterogeneity. Consider example 1.2 from the introduction. In this case, all alternatives are ex ante identical, so there is no heterogeneity in item quality. This in turn means that the default choice probability is the same in any choice set with the same cardinality. Thus, collecting (for example) data from multiple sets of size two and applying the Min bound does not help. However, as the example demonstrates, the RUM bound based on observations from choice sets of size two and three does put further constraints on the data. This is because the assumption that randomness is driven by preference heterogeneity puts extra structure on the allowed choice probabilities. Thus the RUM-based definition of choice overload gives potentially tighter bounds with which to identify choice overload. [^10] Note that, without preference heterogeneity, the RUM bound will equal the Min bounds. If all subjects have the same preferences, then in all small choice sets the default choice probability must be either zero or one. Thus, the Min bound must either be zero, if in some choice set the default is never chosen, or one, if the default is always chosen in every choice set. In both cases the RUM bound would be the same as the Min bound.
A disadvantage of RUM bound is that it requires data on $\mathcal{D}/X$ to be consistent with RUM. There are two possible issues with this. First, it could be that the population choice distributions on on $\mathcal{D}/X$ are inconsistent with RUM; in that case, $p_{D}^{RUM}(X)$ is not well-defined. Second, it could be that rationalizable population distributions generate non-rationalizable finite sample frequencies; in that case, one could define a feasible version of $p_{d}^{RUM}(X)$, and we will do so in Section 4.3.
It remains to clarify how we can test stochastic rationalizability of data. In its essence, this question has been resolved in McFadden and Richter (1991). Because we adapt their approach to an environment in which rationalizability is not characterized by standard revealed preference axioms (although they mention the possibility of this generalization) and to keep the treatment self-contained, we will briefly explain the answer. The basic insight is that probabilistic choice data can be rationalized if, and only if, it can be expressed as convex combination of data that would be produced by deterministic utility maximizers. [^11] As is often the case with seminal results, with ample hindsight the intuition informs a much easier than the original proof (Stoye, 2019).
To make this precise, construct a matrix $\bm{A}$ s.t. each row of $\bm{A}$ corresponds to a given alternative within a particular choice set (i.e., each alternative appears once for every choice set containing it), and each rationalizable deterministic choice pattern corresponds to exactly one column of $\bm{A}$. For example, consider a data set consisting of choices from $\left\{a_{1},d\right\}$, $\left\{a_{1},a_{2},d\right\}$ and $\left\{a_{1},a_{2},a_{3},d\right\}$. The first row of $\bm{A}$ indicates the choice of $a_{1}$ from $\{a_{1},d\}$, the second row the choice of $d$ from $\{a_{1},d\}$ and so on. One column of $\bm{A}$ (namely, the first one in 3.1 below) then represents the choices of someone who picked $a_{1}$ from all three choice sets, which can be rationalized as the choices of a decision maker who prefers $a_{1}$ over all other alternatives in $X$. The remaining columns of $\bm{A}$ represent all other rationalizable choice patterns. Some book-keeping reveals that in this example, one has: [^12] The size of $\bm{A}$ increases extremely rapidly with the complexity of data; see Kitamura and Stoye (2018) for numerical examples.
$\begin{array}[]{c}a_{1}|\left\{a_{1},d\right\}\\ d|\left\{a_{1},d\right\}\\ a_{1}|\left\{a_{1},a_{2},d\right\}\\ a_{2}|\left\{a_{1},a_{2},d\right\}\\ d|\left\{a_{1},a_{2},d\right\}\\ a_{1}|\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{2}|\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{3}|\left\{a_{1},a_{2},a_{3},d\right\}\\ d|\left\{a_{1},a_{2},a_{3},d\right\}\end{array}\left\{\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&0&0&0&1&1&1&1\\ 1&1&0&0&0&0&0&0\\ 0&0&1&1&0&0&1&1\\ 0&0&0&0&1&1&0&0\\ 1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&1&0\\ 0&1&0&1&0&1&0&1\\ 0&0&0&0&1&0&0&0\end{array}\right\}=\bm{A}.$ (3.1)
The core insight of McFadden and Richter (1991) is summarized in the following theorem.
Theorem 3.1.
Let $\pi$ be the observed choice probabilities associated with each row of the matrix $\bm{A}$ derived from some stochastic choice data set. That data set is rationalizable by RUM if and only if there exists a vector $\nu\in\Delta^{H-1}$ (the $(H-1)$-dimensional unit simplex, where $H$ is the number of columns of $\bm{A}$) such that
$\bm{A}\nu=\pi$
In order to adapt this approach to our idealized data set, we need to deal with the fact that we only observe whether or not the default was chosen. To do so, we premultiply the matrix $\bm{A}$ with a matrix $\bm{B}$ that merges events we do not separately observe – i.e. the choice of different non-default options. In the above example, we would have
$\bm{B}=\left\{\begin{array}[]{ccccccccc}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&1&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right\}$
A corollary to Theorem 3.1 then characterizes $p_{d}^{RUM}(X)$:
Corollary 3.1.
Let $\pi$ be the observed choice probabilities associated with each row of the matrix $\bm{BA}$ derived from a data set $\{\mathcal{D},p\}$ which reports only the probability of default choice. That data set is rationalizable by RUM if and only if their exists a probability vector $\nu$ such that
$\bm{BA}\nu=\pi.$
Further, let $\bm{a}$ be the row of $\bm{A}$ that corresponds to default choice from $X$, then
$p_{d}^{RUM}(X)=\max_{\nu\geqq 0}\{\bm{a}\nu\}\text{ s.t. }\bm{BA}\nu=\pi.$ (3.2)
For intuition, observe that in (3.1), the vector $\bm{a}$ is the last row of $\bm{A}$. Equivalently, it is the indicator of the choice type being the one whose choice is always the default (this is the only type who chooses the default from $X$). The bound simply maximizes the probability corresponding to this type, subject to overall data being stochastically rationalizable.
In practice, $p_{d}^{RUM}(X)$ as stated is well-defined only if $\pi$ is stochastically rationalizable. Since the set of rationalizable probability vectors is “small” (we will elaborate on this in Section 4.3), empirical choice frequencies will typically –and do in our data– fail this. In that case, a feasible version of the bound can be computed by substituting a constrained estimator of $\pi$. This will be illustrated later.
We finally note that the above machinery opens the door to using this testing approach on other, for instance non-RUM, models. Nothing in the above development forces $\bm{A}$ to contain columns that correspond exactly to conventionally rationalizable behaviors. By adding (removing) columns of $\bm{A}$, one can test less (more) restrictive models. Indeed, in our empirical application we will use this approach to test:
  • (i)
    The RUM as just explained.
  • (ii)
    A model which admits all the rationalizable deterministic choice patterns of RUM but also allows for choosing $d$ from choice set $X$ (regardless of behavior on smaller sets).
  • (iii)
    The same model as (ii), except that a choice type may also switch to choosing $d$ for all choice sets of cardinality $3$. (Types that choose $d$ from all sets of cardinality $3$ must also choose $d$ from choice set $X$.)
Models (ii) and (iii) capture ‘choice overloaded’ behavior in that the only violations of standard rationality they allow are switches to the default when the choice set expands. Moving from (i) to (iii) enlarges $\bm{A}$, and therefore increases computational cost, but does not add conceptual difficulties. Since the resulting models are nested (listed in increasing order of generality), one can then ask what is the least permissive model that is not rejected in the data. This will allow us to unpack what sort of choice overloaded behavior, if any, could have generated our data; for reasons that will become clear, this line of analysis will be central to our results. As a preview, model (iii) will be the only one not rejected in our data. This analysis is in Section 4.3, where we also argue that model (iii) is itself a restrictive model.
We note that a further generalization which captures choice overload would be to allow switches to the default not as a function of set cardinality (as in (ii) and (iii)), but along any distinct path of choice set expansion. Dean (2008) considers a similar model. We do not consider this generalization due to computational complexity (the corresponding $\bm{A}$-matrix is very large) and because we know it would not be rejected in our data given it nests (iii).
3.2 Econometric Tests
We next explain some testing strategies that connect the above ideas to recent advances in econometrics. To this purpose, we consider samples that were generated by randomly drawing individuals and then exposing each individual to an (i.i.d. randomly generated) selection of choice problems, i.e. subsets of $\mathcal{D}$ [^13] This mirrors our empirical design, which was partly chosen because, unlike stratified sampling or mean-reverting coins, it is easy to bootstrap. In particular, data may contain choices from the same individuals in different choice sets, as is the case in our application. Note also that in this section, the term “data” is used in the statistical sense, i.e. referring to a finite sample from an underlying population.
We begin with two tests of the Min bound. We will estimate choice probabilities $p_{d}(\cdot)$ by the analogous sample frequencies, which can be informally defined as
$\hat{p}_{d}(A)\equiv\frac{\sum\bm{1}\{\text{choice problem is }A,\text{ choicemade is }d\}}{\sum\bm{1}\{\text{choice problem is }A\}},$ (3.3)
where $\bm{1}\{\cdot\}$ is the indicator function and the sums are taken over all data points. For all but the finite-sample test that we present first, any estimator of $p_{d}(\cdot)$ whose asymptotic distribution is normal or approximated by the simple nonparametric bootstrap would suffice. Sample averages have both properties as long as they are not close to degenerate; by direct inspection, this condition is easily met in our data.
3.2.1 A Finite-Sample Test of the Min Bound
Testing the Min bound amounts to testing whether
$p_{d}(X)\leq\min_{A\in\mathcal{D}\setminus X}p_{d}(A)$
$\Longleftrightarrow$ $p_{d}(X)\leq p_{d}(A),\forall A\in\mathcal{D}\setminus X.$
The second expression clarifies that this is a joint test of potentially many hypotheses. Consider first testing any one of these, i.e. testing whether $p_{d}(X)\leq p_{d}(A)$ for a specific $A\subset X$. To this purpose, define $\hat{p}_{d,A}(X)$ in exact analogy to (3.3) but dropping observations from subjects who also saw choice problem $A$. In words, we estimate $p_{d}(X)$ by its simple empirical analog among subjects who did not see $A$. As a result, $\hat{p}_{d}(A)$ and $\hat{p}_{d,A}(X)$ estimate binomial proportions in two mutually exclusive samples. We will therefore apply Fisher’s (1992) exact test for binomial proportions to $H_{0}:p_{d}(X)\leq p_{d}(A)$ for any given $A$.
Of course, we need to account for the fact that we will conduct many such tests once ($78$ in our application) and cannot assume independence. Our first approach is Bonferroni adjustment, that is, all p-values are multiplied by $78$. An advantage of this approach is that it ensures finite-sample (as opposed to asymptotic) size control. However, its power is limited through three channels: The estimators $\hat{p}_{d,A}(\cdot)$ discard data; Fisher’s exact test is in general conservative due to integer issues; and Bonferroni adjustment is conservative. In practice, with the sample sizes that we generated for our empirical application, we expect only the last channel to have an appreciable effect.
3.2.2 An Asymptotic Test of the Min Bound
The finite sample test adjusts for the fact that, in principle, many tests are conducted simultaneously. A common concern with such adjustments is that, if the results of some of these tests appear obvious, one might needlessly lose power. Indeed, consider the visualization of our data in Figure 3: The default probabilities in the right-hand cluster are obviously much higher than in the grand choice set (corresponding to the dashed line). Can we restrict attention to only those inequality conditions that might reasonably bind?
This question has received considerable attention in the econometric literature on moment inequalities and is by now well understood. We implement a method that can be seen as special case of Andrews and Soares (2010) and also of Chernozhukov et al. (2013), both of whom establish its validity under rather general conditions.
The method can in principle use many test statistics; for concreteness, set
$t=\hat{p}_{d}(X)-\min_{A\in\mathcal{D}\setminus X}\hat{p}_{d}(A).$
The test will reject if $t$ is too large. The catch is that the distribution of $t$, and therefore the appropriate critical value, depends on the nuisance parameter $(p_{d}(X)-p_{d}(A))_{A\in\mathcal{D}\setminus X}$. This parameter cannot be pre-estimated with sufficient accuracy [^14] Technically, it enters the asymptotic distribution scaled by $\sqrt{n}$. Our exposition is slightly informal because the issue is well understood in the literature; see also Canay and Shaikh (2017, section 4) for a survey. and so we must conservatively approximate it. This is done in three steps:
  1. 1.
    Use the simple nonparametric bootstrap to approximate the distribution of
    $(\hat{p}_{d}(A)-p_{d}(A))_{A\in\mathcal{D}}$
    by the (bootstrap) distribution of
    $(\tilde{p}_{d}(A))_{A\in\mathcal{D}}\equiv(\hat{p}_{d}^{*}(A)-\hat{p}_{d}(A))_{A\in\mathcal{D}},$
    where $\hat{p}_{d}^{*}(\cdot)$ denotes the bootstrap analog of $\hat{p}_{d}(\cdot)$. This bootstrap will be clustered by individual, i.e. we (i.i.d. uniformly with replacement) resample individuals and use all responses from a given resampled individual; this ensures that correlation patterns in $(\hat{\pi}-\pi)$ due to eliciting several responses per individual are captured.
  2. 2.
    Use a pre-test with size converging to 00, e.g. $\alpha_{n}=1/\log(n)$, and discard from consideration any sets $A$ s.t. the null hypothesis $H_{0}:p_{d}(X)\geq p_{d}(A)$ is rejected at significance level $\alpha_{n}$. Let $\mathcal{D}^{*}$ denote the set of choice problems that are retained in this pre-test.
  3. 3.
    The critical value of our test is the appropriate quantile of the recentered bootstrap test statistic
    $t^{*}\equiv\tilde{p}(X)-\min_{A\in\mathcal{D}^{*}\setminus X}\tilde{p}(A).$
This procedure reflects two important ideas from the moment inequalities literature. First, the bootstrap population of data must be on the null hypothesis, which necessitates a recentering. In our case, the least favorable and therefore relevant instance of the null hypothesis is that all relevant probabilities are equal. Since the test statistic is location invariant, for concise notation and implementation we recenter them to 00. This is reflected in the definition of $\tilde{p}_{d}(\cdot)$. Second, the test may be extremely conservative if we accordingly recenter all $78$ estimators. Therefore, we pre-screen choice items whose default probability is likely to much exceed $p_{d}(X)$. This is a special case of Generalized Moment Selection (Andrews and Soares, 2010). In particular, the size of the pre-test must go to 00 in order for us to claim that we will asymptotically select all binding constraints. [^15] Our depiction of the method is simplified by picking a specific test statistic and also filling in specific values for several tuning parameters. For example, the sum $\sum_{A\in\mathcal{D}^{*}\setminus X}\max\{\hat{p}_{d}(X)-\hat{p}_{d}(A),0\}$ of constraint violations might yield a test statistic that is more powerful against small but uniform violations. Also, rather than letting the size of a pre-test vanish, one could use Bonferroni correction to “spend” some “coverage budget” on the pre-test (Andrews and Barwick, 2012; Romano et al., 2014). Again, we omit these generalizations for conciseness because they are known in the literature. The test that we implement is as stated, and the rejection in our empirical application is so resounding that none of these choices would plausibly have impacted it.
3.2.3 An Asymptotic Test of RUM and its Generalizations
Statistical testing of random utility models is due to Kitamura and Stoye (2018), with important computational improvement by Smeulders et al. (2020). It has seen application to observational data (Deb et al., 2022) as well as lab experiments (Aguiar et al., 2022). Because we will slightly adapt the approach to our setting, and also for self-contained exposition, we next briefly explain it. We also remind the reader that RUM is here meant in a somewhat loose sense: It is not essential that the population is modelled as mixing a finite number of “admissible” types encoded in the columns of $\bm{A}$, not that admissibility coincides with standard economic rationality; hence, we can use the machinery to test nonstandard and, in particular, nested models.
The Hypothesis Test
The null hypothesis
$H_{0}:\bm{BA}\nu=\pi,\exists\nu\in\Delta^{H-1}$
can equivalently be written as
$H_{0}:$ $\bm{BA}\nu=\pi,\exists\nu\geq 0$
$\Longleftrightarrow$ $\pi\in\mathcal{C}\equiv\{\bm{BA}\nu:\nu\geq 0\}$
$\Longleftrightarrow$ $\min_{\nu\geq 0}\{(\pi-\bm{BA}\nu)^{\prime}\Omega(\pi-\bm{BA}\nu)\}=0,$
where $\Omega$ is an arbitrary positive definite (and in practice diagonal) weighting matrix. Here, the first step is true because a vector $\nu$ can fulfil $\bm{BA}\nu=\pi$ only if its elements sum to $1$, thus we can relax the constraint on $\nu$. The last expression states that the residuals from projecting $\pi$ onto the cone $\mathcal{C}$ of rationalizable probabilities must equal 00 and suggests the scaled norm of the corresponding sample residuals as natural test statistic. Noting the similarity to specification tests in multiple equation models (Sargan, 1958; Hansen, 1982), call this statistic
$J_{n}\equiv n\min_{\nu\geq 0}\{(\hat{\pi}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}-\bm{BA}\nu)\},$
where $\hat{\pi}$ stacks estimated choice probabilities $(\hat{p}_{d}(A),1-\hat{p}_{d}(A))$ in an order corresponding to $\pi$ and $n$ is sample size.
The asymptotic distribution of $J_{n}$ is delicate to estimate because it depends discontinuously on where on $\mathcal{C}$ the true $\pi$ is. Valid inference therefore relies on a bootstrap procedure that not only recenters the empirical distribution to be on $H_{0}$ (as in the previous test) but that also shrinks $H_{0}$ in a locally unfavorable direction. Heuristically, the idea is once again similar to Generalized Moment Selection in moment inequalities, but the details differ because, while $\mathcal{C}$ is a finite polytope, we can only characterize it in terms of its vertices (the columns of $\bm{A}$) as opposed to its faces (which would correspond to linear inequality constraints and allow for Generalized Moment Selection if we had them).
Operationally, we recenter the bootstrap distribution of data at the projection onto the tightened cone
$\mathcal{C}_{\tau_{N}}\equiv\{\bm{BA}\nu:\nu\geq\bm{1}\cdot\tau_{n}/H\},$
where $\tau_{n}$ is a tuning parameter that we set equal to $\sqrt{\log(\underline{n})/\underline{n}}$ with $\underline{n}$ being expected sample cell size for any but the universal choice problem; $\bm{1}$ is a vector of $1$’s. [^16] Simple algebra reveals that $\underline{n}=\tfrac{2qn}{k(k+1)}$, where $k$ is the number of nondefault items in $X$ and $q$ is the number of choice problems other than the grand problem faced by each subject. Recall also that $H$ is the length of $\nu$, thus division by $H$ ensures that the above constraint is scaled by $\tau_{n}$ and not by the testing problem’s complexity. The essential feature here is that $\tau_{n}$ vanishes but slowly compared to estimation uncertainty in $\hat{\pi}$. Then the bootstrap analog of $J_{n}$ is defined as
$J^{*}_{n}$$\equiv$ $\min_{\nu\geq\bm{1}\cdot\tau_{n}/H}\{(\hat{\pi}_{\tau_{n}}^{*}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}_{\tau_{n}}^{*}-\bm{BA}\nu)\}$ (3.4)
$\hat{\pi}_{\tau_{n}}^{*}$$\equiv$ $\hat{\pi}^{*}+\hat{\eta}_{\tau_{n}}-\hat{\pi}$
$\hat{\eta}_{\tau_{n}}$$\equiv$ $\arg\min_{\nu\geq\bm{1}\cdot\tau_{n}/H}\{(\hat{\pi}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}-\bm{BA}\nu)\}$ (3.5)
and $\hat{\pi}^{*}$ is a simple nonparametric (clustered, as explained earlier) bootstrap analog of $\hat{\pi}$.
Validity of the resulting test under reasonable assumptions is established in Kitamura and Stoye (2018). We go beyond a completely straightforward implementation of their test because we do not weight questions equally. This possibility is anticipated in Kitamura and Stoye (2018) because they require only a diagonal weighting matrix $\Omega$ but has not, to our knowledge, been implemented before. It is motivated by the fact that, in our data, choice probabilities pertaining to the universal choice set $X$ will be estimated from a much larger sample cell, and therefore much more precisely, than others. Specifically, both the expected and also realized average relative cell size of the last versus other questions are determined by the experimental design, and it is this number that we will use, i.e. the weighting matrix will be
$\Omega=\left(\begin{array}[]{rrrrr}1&\cdots&0&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ 0&\cdots&1&0&0\\ 0&\cdots&0&w&0\\ 0&\cdots&0&0&w\end{array}\right),$
where $w$ is easy to compute and takes a value of $w\approx 9$ in our experimental design. [^17] Indeed, $w=\tfrac{k(k+1)}{2q}$, with $(k,q)$ as in the preceding footnote. We do not weight questions by realized sample cell sizes, and we also do not estimate cell-specific variances by the binomial variance formula, in order to avoid data dependent weighting. In our application, these modifications would have minimal effect.
We close by observing that in general, this test can be expensive to compute. The experimental design that we settled on, partly to ensure reasonable sample cell sizes, is small enough so that this is not the case. However, in preparation, we also implemented an adaptation of the computational improvements in Smeulders et al. (2020). For the benefit of future users, these details are laid out in Appendix B.
4 An Application to Experimental Data
We next describe an application of the above methods to an experimental data set.
4.1 Experimental Design
The aim of the experiment is to collect data of the type needed to implement tests for the Min and RUM bounds. This means we require repeated observations of choices from a grand set of alternatives and a number of subsets, with all sets containing a default option. Moreover, based on the findings of Chernev et al. (2015), we want the choice problems to be non-trivial to increase the probability of finding choice overload. To this end, we ask subjects to make choices between amounts of experimental points expressed as sums, as in Caplin et al. (2011).
Specifically, each non-default option is expressed as a sum of four numbers between zero and ten written in text. The value of choosing an option in experimental points is the value of the sum, and one experimental point is worth 50 cents. There are 12 non-default options in the grand set $X$. Following Caplin et al. (2011), we generated these as follows: First, we drew the value of an alternative was drawn from an exponential distribution truncated at 10 points with $\lambda=0.25$. Next, individual terms of sums were chosen stochastically ensuring neither the first nor the maximal summand was correlated with the total value of the option.
Meanwhile, a default option was available in each choice set. This option provided 7 points and was the only option expressed as a single number. The default option also always appeared at the top of the screen, was pre-selected (i.e. the subject had to actively click on another option if they wanted to choose it) and the simplest alternative (as measured by the number of summands in the expression). Thus it fulfilled many criteria that have been proposed to define the “default” in choice overload situations (see for example Iyengar and Kamenica (2010)). Of the non-default options, 9 yielded a number of points strictly lower than the default, 1 yielded the same number of points as the default, and 2 yielded strictly more points. Figure 1 shows an example choice screen with the default and one other option. Figure 2 shows an example choice screen for the grand set.
Figure 1: A choice problem comparing one lottery to the default.
Figure 2: A choice problem comparing the grand set to the default.
The collection of choice sets used in the experiment consisted of the grand set (12 alternatives plus the default), and all 1 and 2 alternative subsets (along with the default), for a total of 78 smaller choice sets. Based on the prior literature, we initially believed that 3 item sets are unlikely to trigger choice overload, while 13 item sets are likely to do so, if such an effect is present. Each subject was presented with 9 randomly selected small sets and the grand set, with the order of choice questions and the order of non-default options in each choice question randomized. One question was randomly selected for payment. Subjects in addition received a \$1 participation fee. A complete list of the choice alternatives and choice sets can be seen in Appendix A.2.181818The experiment was approved by the Institutional Review Boards of Columbia and Cornell Universities.
Note that we have selected a choice environment that should generate a good deal of heterogeneity in item quality, but in which there should be little preference heterogeneity. Based on the discussion of Section 3.1 we would anticipate the Min bound to be effective in identifying choice overload, whle the RUM bound may not necessarily provide further advantages.
Subject Recruitment. The experiment was run on Amazon’s Mechanical Turk (MT) platform, a digital marketplace for work. This platform was chosen to make it easy to get data from a large number of subjects, each of whom needed to answer only a limited number of questions. “Requesters” post Human Intelligence Tasks, or “HITs,” which are usually simple, repetitive jobs that typically pay small sums for each completed task. Workers on MT view descriptions of the HITs, decide which to accept, and complete those HITs over the internet. In our case, subjects who accepted the HIT followed a link to an external webpage, where they completed the experiment. Upon completion they were given a randomly generated code, which was used to pay them the appropriate amount given their choices in the experiment.
Instead of posting HITs and recruiting subjects directly on MTurk, we used an online platform called CloudResearch. CloudResearch posts our experiment as a HIT on MTurk and recruits the subjects. All subjects recruited through CloudResearch’s platform are first screened through CloudResearch’s Sentry data validation system. Sentry is a short (1 minute) pre-experiment survey meant to ensure subjects are attentive, engaged, and ready to participate in the experiment. People who fail Sentry are routed away from the experiment. These pre-study validation surveys have been shown to be effective in increasing data quality (Chandler et al., 2019; Litman and Robinson, 2020).
We recruited 2000 subjects for the experiment. In addition to CloudResearch’s screening, we restricted recruitment to MTurk workers who had completed over 1000 HITs and had an approval rating of over 97%. Following the instructions subjects had to answer two quiz questions to ensure the understood the instruction. 1833 subjects passed the quiz and completed the experiment, but one of these subjects had a browser related error which made their data unusable, leaving us with 1832 subjects’ data.
The resulting data provides 1832 choices from the grand set. We have at least 185 observations of choices from each of the small sets, with variation in sample size due to the random selection of choice questions.
4.2 Data Overview
Figure 3 presents a histogram showing the distribution of choice sets by the probability with which the default option was chosen from sets other than $X$. The distribution is bimodal. The choice sets with low probability of default choice are precisely the ones in which the best option yields more points than the default. The probability of default choice in $X$ is indicated by the dotted vertical line. It falls somewhat above the default choice probability of the lower group, but below that of the higher group. Appendix A.2 lists the default choice probabilities for each choice set.
Figure 3: Histogram showing distribution of choice sets by probability that default option was chosen. Dotted line shows the fraction of default choice in the grand set $X$.
4.3 Analysis
We next ask whether the data we collected are suggestive of choice overload by conducting tests in roughly increasing order of presumed power.
First, we establish that existing approaches would be unlikely to identify choice overload in our data set. The empirical frequency of default choice in the grand set, $\hat{p}_{d}(X)$, was $22.3\%$. The analogous empirical frequency across all subsets was $70.7\%$ and therefore considerably higher. A simple comparison of means would not reveal choice overload, but also masks extreme heterogeneity across choice items. A random selection of a single smaller set would also be unlikely to reveal choice overload. $23$ of $78$ small choice sets had default choice frequency below $22.3\%$, only $15$ of which significantly so at the 5% level (using Fisher’s exact test on disjoint samples but without controlling for multiple comparisons). Thus, if a researcher were to randomly select a small choice set to compare to the grand set, they would only find evidence for choice overload 19% of the time.
In a next step, we use the finite sample test to ask whether any of those $23$ frequencies is significantly below the analogous frequency in the grand choice set, taking into account sampling uncertainty. Strikingly, the answer is yes: The p-value against the null hypothesis that the grand choice default frequency is lowest is .00003.00003.00003.00003, and $5$ choice items are significantly below it at the $5\%$-level, where the p-values and significance levels just reported reflect Bonferroni adjustment across $78$ tests. [^19] The choice sets alluded to are numbers $(9,11,20,32,74)$ in order of appearance in Table 2. Similarly, the asymptotic test of the Min bound yields a p-value that we could not distinguish from 00 in $B=10000$ Monte Carlo simulations. Hence we find strong evidence of choice overload using this approach.
We next apply the test explained in Section 3.2.3 to see if our data is consistent with RUM. At a Monte Carlo replication size of $B=10000$, the p-value against this null hypothesis is equals 00 as well. This extremely strong rejection comes with a caveat: The p-value against all but the grand choice set equals .005.005.005.005. Therefore, “the data are consistent with RUM except that default choice from $X$ is too frequent” is not an appropriate description of our findings. However, this leaves open the possibility that choice overload had an effect already in some of the smaller choice sets.
[…]
Second, to further quantify the sense in which these models fit the data, we compute the largest sample proportion of subjects such that individually rational behavior by these subjects would be compatible with empirical choice frequencies. This proportion is defined by the linear program
$\max\bm{1}^{\prime}\nu\text{ s.t. }\bm{A}\nu\leq\hat{\pi}$
and equals $1$ if, and only if, sample proportions are rationalizable in terms of rational matrix $\bm{A}$. This fraction is .866.866.866.866 for the RUM, i.e. model (i) above, increases to .877.877.877.877 for model (ii), and equals .915.915.915.915 for model (iii).
Third, one may worry that model (iii) is just not very restrictive, while maybe models (i) and (ii) are. This cannot be literally true because our test statistics are positive in all three tests, hence empirical choice frequencies are not rationalizable under any model. But it could certainly be “morally true,” notably if all possible data sets are “close” to the model. In order to shed some light on this, we conducted some analyses inspired by Bronars (1987), Selten (1991), and Beatty and Crawford (2011). Specifically, for models (i)-(iii) we calculated the expected mean square error (MSE) when matching data that were generated from the uniform distribution on $[0,1]^{79}$ (i.e. choice probabilities drawn uniformly at random). The resulting values are 0.25 for model (i), 0.23 for (ii) and 0.21 for (iii). These numbers are all close to each other, and far above the MSE when these models are applied to the data (which range from 0.03 to 0.01), showing that even our most permissive model places significant restrictions on the data.
Finally, the data can be used to illustrate the difference between the Min bound, $p_{d}^{min}(X)$, and the RUM-based bound, $p_{d}^{RUM}(X)$. The empirical choice frequencies do not imply a well-defined $p_{d}^{RUM}(X)$ because they are not stochastically rationalizable even if choice set $X$ is ignored. However, one can easily compute the vector $\hat{\eta}$ of choice probabilities that are closest to the empirical ones while being rationalizable; in fact, this computation is a by-product of the statistical tests. This allows us to compute a feasible analog of $p_{d}^{RUM}(X)$. Its value in our data is $9.8\%$. To compare apples to apples, we report that the same $\hat{\eta}$ implies a Min bound of $p_{d}^{min}(X)=11.4\%$. Using the full implications of RUM is potentially much more informative than just testing monotonicity.
5 Conclusion
In this paper, we have argued that existing tests for choice overload are underpowered, and that this may explain the confusing and contradictory picture that emerges from the current literature. We proposed that, by collecting more data and making use of theory, one can design better tests. We have demonstrated that, in a novel data set, our approach finds choice overload. Indeed, our new data speak so loudly that with hindsight, our methodological innovations would not have been necessary to detect choice overload in our own data set. We believe that the innovations are of interest nonetheless, and we also note that such things are bound to occur if –as we did– one genuinely designs the empirical strategy before going into the field.
One unexpected finding from our experiment is that choice appears to violate the RUM even in smaller choice sets. This is perhaps surprising given the relative simplicity of the choice items. Understanding the cause of these violations strikes us as an interesting avenue for future research. For our purposes, it suggests that RUM may not be an appropriate benchmark against which to look for choice overload, at least in this experimental task.
Appendix A Supplementary Appendix: Experimental Details
Here we present the experiment instructions, quiz, and details about the choice objects.
A.1 Instructions and Quiz
Subjects were first shown this instructions screen before proceeding to the quiz (below). Subjects were given two attempts at the quiz.
[Uncaptioned image] [Uncaptioned image]
A.2 Choice Alternatives with Summary Data
Table 1 gives full list of the choice objects with their values in experimental points.
Table 1: List of Options with Values
ID Option Value
0 seven 7
1 eight minus seven plus eight minus nine 0
2 eight minus one plus two minus seven 2
3 seven minus seven minus one plus two 1
4 six minus ten plus one plus five 2
5 seven minus eight plus nine minus two 6
6 five plus eight plus zero minus nine 4
7 nine minus eight plus two plus four 7
8 nine minus eight minus ten plus ten 1
9 four minus two minus one plus seven 8
10 two plus six minus four minus three 1
11 three plus zero plus nine minus two 10
12 six minus ten plus seven minus two 1
Table 2 shows how often each choice set was shown to subjects and how often the default was chosen from it. Choice sets are given by the alternatives they contain: $[0,1,2]$ represents the choice set with the default (0) and options with IDs 1 and 2 from Table 1.
Choice Set # Choices # Default Choice Set # Choices # Default
$[0,1]$ 204 199 $[0,3,10]$ 199 190
$[0,2]$ 193 188 $[0,3,11]$ 200 38
$[0,3]$ 218 210 $[0,3,12]$ 231 219
$[0,4]$ 232 225 $[0,4,5]$ 219 190
$[0,5]$ 227 214 $[0,4,6]$ 215 200
$[0,6]$ 230 226 $[0,4,7]$ 213 191
$[0,7]$ 221 212 $[0,4,8]$ 216 187
$[0,8]$ 229 212 $[0,4,9]$ 193 35
$[0,9]$ 201 18 $[0,4,10]$ 219 204
$[0,10]$ 199 190 $[0,4,11]$ 210 28
$[0,11]$ 194 20 $[0,4,12]$ 197 186
$[0,12]$ 195 184 $[0,5,6]$ 224 200
$[0,1,2]$ 209 203 $[0,5,7]$ 209 183
$[0,1,3]$ 225 217 $[0,5,8]$ 210 186
$[0,1,4]$ 185 175 $[0,5,9]$ 224 45
$[0,1,5]$ 204 190 $[0,5,10]$ 199 189
$[0,1,6]$ 208 199 $[0,5,11]$ 214 36
$[0,1,7]$ 203 188 $[0,5,12]$ 213 199
$[0,1,8]$ 225 203 $[0,6,7]$ 205 189
$[0,1,9]$ 211 24 $[0,6,8]$ 200 178
$[0,1,10]$ 218 215 $[0,6,9]$ 218 44
$[0,1,11]$ 223 30 $[0,6,10]$ 209 204
$[0,1,12]$ 235 219 $[0,6,11]$ 223 31
$[0,2,3]$ 229 219 $[0,6,12]$ 221 210
$[0,2,4]$ 213 202 $[0,7,8]$ 223 202
$[0,2,5]$ 218 202 $[0,7,9]$ 206 36
$[0,2,6]$ 215 208 $[0,7,10]$ 199 182
$[0,2,7]$ 250 231 $[0,7,11]$ 205 30
$[0,2,8]$ 193 178 $[0,7,12]$ 221 205
$[0,2,9]$ 207 36 $[0,8,9]$ 226 32
$[0,2,10]$ 192 185 $[0,8,10]$ 205 182
$[0,2,11]$ 194 24 $[0,8,11]$ 222 33
$[0,2,12]$ 192 182 $[0,8,12]$ 221 204
$[0,3,4]$ 191 182 $[0,9,10]$ 192 31
$[0,3,5]$ 218 203 $[0,9,11]$ 202 23
$[0,3,6]$ 198 194 $[0,9,12]$ 223 42
$[0,3,7]$ 205 192 $[0,10,11]$ 219 33
$[0,3,8]$ 215 194 $[0,10,12]$ 193 187
$[0,3,9]$ 207 44 $[0,11,12]$ 224 36
$[0,\ldots,12]$ 1832 409
Table 2: Choice Set and Default Choice Frequencies
Appendix B Implementing recent computational innovations
We implemented the computational procedure in Smeulders et al. (2020). To our knowledge, this is the first such implementation beyond their own illustrative example. The implementation and some not entirely obvious modifications are described next.
The procedure is motivated by the fact that computation of the matrix $\bm{A}$ and also computation (3.4) is hard, and yet the latter needs to be repeated many times. It exploits that, because $\bm{BA}$ has many more columns than rows, there always exists a sparse (in the loose sense of having relatively few nonzero entries) $\arg\max$ to problem (3.4). We will avoid solving (3.4) as stated, or ever computing $\bm{BA}$ (though the latter is feasible here), by guessing the nonzero entries. Formally, this goes as follows. (We drop $N$ for ease of notation.):
  1. 1.
    Initialize the matrix $\tilde{\bm{BA}}$ by constructing relatively few columns of $\bm{BA}$.
  2. 2.
    Compute
    $\tilde{J}\equiv\min_{\nu\geq 0}\{(\hat{\pi}-\tilde{\bm{BA}}\nu)^{\prime}\Omega(\hat{\pi}-\tilde{\bm{BA}}\nu)\}.$
    Let $\tilde{\eta}\equiv\tilde{\bm{BA}}\tilde{\nu}$, where $\tilde{\nu}$ solves this problem. (While $\tilde{\nu}$ may not be unique, $\tilde{\eta}$ is.)
  3. 3.
    Maximize $(\hat{\pi}-\tilde{\eta})^{\prime}\Omega(a-\tilde{\eta})$ subject to the constraint that $a$ is a column of $\bm{BA}$. This is called the “pricing problem.” Its constraint must be expressed in an application specific, computable way, and we do so below.
  4. 4.
    If the value of the problem just solved is positive, append column $a$ to $\tilde{\bm{BA}}$. Repeat until the value of the problem is nonpositive or another convergence criterion is met.
The basic idea is that, as long as the deficient matrix $\tilde{\bm{BA}}$ does not contain all columns that receive positive weight in one solution to the original problem, the value of the simplified problem can be improved by appending such a column. But a column improves this value iff the supporting hyperplane separating the current feasible set from $\hat{\pi}$ does not separate the new column from $\hat{\pi}$. The program in step 3 simply checks this. (We solve it but in principle, it suffices to sign its value.) If the solution is sparse, it will be found while only generating a fraction of all possible columns of $\bm{BA}$.
Our implementation is again not completely off the shelf. Modifications are as follows:
  • (i)
    We take account of the weighting matrix $\Omega$ not being the identity matrix. This is already reflected in expressions above.
  • (ii)
    The requirement that the vector $a$ be a possible column of $\bm{BA}$ can be expressed by writing the pricing problem as follows. To enforce that $a$ is binary and any two entries corresponding to the same choice problem sum to $1$, parameterize it in terms of a vector $\rho$ that only collects indicators of active choice. Then $a=d+D\rho$, where
    $d$ $=$ $\left(\begin{array}[]{c}0\\ 1\\ 0\\ \vdots\\ 1\end{array}\right),~{}~{}~{}D=\left(\begin{array}[]{rrrr}1&0&\dots&0\\ -1&0&\dots&0\\ 0&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&-1\end{array}\right).$
    The objective function of the pricing problem becomes
    $(d+D\rho-\hat{\eta})^{\top}\Omega(\hat{\pi}-\hat{\eta})=(D\rho)^{\top}\Omega(\hat{\pi}-\hat{\eta})+const.$
    Constraints on $\rho$ must reflect that (i) $\bm{0}\leq\rho\leq\bm{1}$; (ii) if choice from one set is active, choice from all supersets thereof is active, (iii) if the default option is chosen from all subsets of a set, then it is chosen from the set as well.
    In sum, the pricing problem can be expressed as the following integer linear program:
    $\max_{\rho\in\{0,1\}^{I}}(D\rho)^{\top}\Omega(\hat{\pi}-\hat{\eta})$
[…]
    $\rho_{i}$$\leq$ $\sum_{j=1,\dots,k:x_{j}\in X_{i}}\rho_{j}.$
  • (iii)
    At first glance, the tightened optimization problem (3.4) has no sparse solution, but Smeulders et al. (2020) remedy this. Heuristically, the vector $\bm{1}\cdot\tau_{N}/H$ can be concentrated out of the problem and a problem with sparse solution remains. A catch is that this requires the initial guess $\tilde{\bm{BA}}$ to have the same dimension as the true $A$ (its column cone cannot be contained in a face of the $\mathcal{C}$). Smeulders et al. (2020) generate columns at random and verify that this constraint is met. This will not work here because only one of possibly millions of choice types makes a default choice on the universal set. Random column generation would be unlikely to discover that type, and so we seed $\tilde{\bm{BA}}$ with the corresponding column, $300$ additional random columns, and verify the rank condition. This is a problem and a fix that is likely to apply to other applications of the method as well.

### v3

@@
 Abstract
-Choice overload – by which larger choice sets are detrimental to a chooser’s wellbeing – is potentially of great importance to the design of economic policy.
+Choice overload – in which larger choice sets are detrimental to a chooser’s well-being – is potentially of great importance in the design of economic policy.
 Yet the current evidence on its prevalence is inconclusive.
@@
 1 Introduction
-The standard model of utility maximization tells us that increasing the set of available options can make the consumer no worse off, and may well improve their welfare.
-Since the pioneering study of Iyengar and Lepper (2000), [^1] See also Reibstein et al.
-(1975).
-a body of work in psychology has called this assumption into question.
-The umbrella term “choice overload” covers a number of phenomena by which larger choice sets seem to make people worse off.
-By now there are a huge number of studies investigating various aspects of choice overload; see Scheibehenne et al.
-(2010) and Chernev et al.
-(2015) for recent reviews and meta analyses.
-Some of the output measures used to identify choice overload are hard to interpret using the classic tools of economic analysis; examples include a reduction in ex-post reported satisfaction or lower confidence that the right choice was made.
-Others fall very much in the realm of choice theory.
-In this paper we focus on the observation that larger choice sets may make people more likely to choose a “default” option.
-Iyengar and Lepper (2000) provide an example of this type of behavior in their famous “jam” study.
-On different days a table was set up in a supermarket displaying a range of jams and offering a money-off coupon if a jam was bought.
-On “limited” choice days, six jams were available; on “extensive” choice days, customers faced 18 additional options.
-Strikingly, people were less likely to buy a jam on extensive choice days (2% vs 12% on limited choice days).
-[^2] These are the unconditional probabilities of buying jam.
-Typically the numbers quoted for this study are the probabilities of buying jam conditional on stopping at the table (3% vs 30%).
-However these numbers are more difficult to interpret as there may be selection into who stops at the table.
-Treating “do not buy” as the default choice option, this violates the Independence of Irrelevant Alternatives (IIA) axiom.
-If choice overload could be well established it would have a number of important implications for economics.
-From a positive perspective, it would clearly violate the standard model of utility maximization.
-Moreover, it is not well explained by some of the more obvious behavioral theories.
-For example, most models of reference dependence assume that, for choices that share the same reference point, the decision maker will behave like a standard utility maximizer, ruling out choice overload type effects (Dean, 2008).
-Models of limited attention can also struggle to replicate the phenomenon, if they cannot explain why an overloaded individual does not simply ignore some of the available options.
-[^3] Though see Cattaneo et al.
-(2020, 2021) for examples of models of consideration sets, and de Lara and Dean (2024) for a model of sequential search that can accommodate choice overload-type behavior.
-This points to the need for models including more exotic elements such as regret (Sarver, 2008), rational contextual inference (Kamenica, 2008; Nocke and Rey, 2021) or decision avoidance (Beattie et al., 1994; Dean, 2008; Gerasimou, 2018).
-From a normative perspective, a key tenet of classic economics is that welfare is weakly improved by increasing the set of options available to the consumer.
-This assumption forms the basis of many policy recommendations, yet choice overload would call it into question.
-Unfortunately, current research into choice overload is inconclusive.
-Some direct replications of previous experiments have failed (Scheibehenne, 2008; Greifeneder et al., 2010).
-One recent meta-analysis concluded that the mean measured choice overload effect is zero (Scheibehenne et al., 2010).
-Another one (Chernev et al., 2015) concludes that whether or not choice overload exists may depend a lot on context.
-The aim of this paper is to argue that existing studies are likely to underestimate the degree of choice overload of this type.
-[^4] Despite the above discussion, from now on we will use ‘choice overload’ to refer specifically to the phenomenon of choosing a default option more often in larger choice sets.
-Almost universally, a data set is said to exhibit choice overload only if the default is more likely to be chosen in a larger choice set than it is in some smaller choice set (see section 2).
-Yet intuitively, under the hypothesis of utility maximization one would expect the default to be chosen much less in larger choice sets, because the decision maker (DM) should be more likely to find something that they prefer to the default.
-This lack of power may potentially explain the inconsistent results from existing tests.
-Two examples may further clarify this point.
+Standard economic reasoning asserts that increasing the set of options cannot make the consumer worse off.
+Yet, starting with the pioneering study of Iyengar and Lepper (2000, see also ()), a growing body of work on choice overload – broadly speaking, the idea that people make worse choices in larger choice sets – has called this assumption into question.
+If true, choice overload would have important normative and positive implications for economics.
+It is inconsistent with utility maximization and many classic behavioral models.
+It also calls into question policy recommendations based on giving consumers as large a range of options as possible from which to choose.
+Unfortunately, current research on choice overload is inconclusive.
+Some direct replications of previous experiments failed (Scheibehenne, 2008; Greifeneder et al., 2010).
+One recent meta-analysis concludes that the mean measured choice overload effect is zero (Scheibehenne et al., 2010), and one (Chernev et al., 2015) concludes that its relevance may heavily depend on context.
+At the very least, there remains debate as to whether choice overload is a widespread and robust phenomenon.
+In this paper, we argue that existing studies are likely to underestimate the extent of choice overload.
+A standard experimental paradigm for identifying overload is to compare the frequency of choosing a default option in small and large choice sets – for example, sticking with the default of not buying a jam in Iyengar and Lepper’s (2000, IL henceforth) classic study.
+[^1] While other measures of overload are used, such as ex post satisfaction and regret, Chernev et al.
+(2015) shows them to be highly correlated.
+A data set is said to exhibit choice overload if and only if the default is more likely to be chosen in a larger choice set than it is in the smaller choice set.
+Yet intuitively, under the hypothesis of utility maximization one would expect the default to be chosen much less in larger choice sets as the number of potentially attractive options grows.
+This lack of power may explain inconclusive results from existing tests.
+Three examples may further clarify this point.
 Example 1.1.
@@
 Example 1.2.
-The following example is even more stylized but avoids distributional assumptions.
-Let the choice universe be $X=\{a,b,c,d\}$, where $d$ stands for default.
-Suppose $a,b,c$ are individually chosen with probability $20\%$ from choice sets $\{a,d\},\{b,d\},\{c,d\}$.
-Suppose also that the probability of active choice is $40\%$ for any of $\{a,b,d\},\{a,c,d\},\{b,c,d\}$.
-Assuming that indifference is not allowed, these probabilities can be reconciled with a Random Utility Model (RUM, precisely defined later); however, they imply that the preferences $a\succ d$, $b\succ d$, and $c\succ d$ occur in nonoverlapping subsets of the population.
-Thus, the active choice probability from $X$ must be at least $60\%$.
-If it is lower, we observe choice overload in the sense that observations cannot be explained in terms of a fully rational population, but they can be explained in terms of a population that is fully rational except that some individuals switch to default choice when encountering the grand choice set.
-This is true even if the active choice probability from $X$ strictly exceeds the active choice probability from any smaller set.
-In this paper, we describe new tests for choice overload.
-We consider a data set based on a finite set of alternatives $X$ that includes a default $d$.
-Observations of choice behavior are made from the grand set $X$, plus some subsets that always contain $d$.
-In each case we assume that the probability that $d$ is chosen is observed.
-We make use of two definitions of choice overload.
-The first is model free: we say that a data set exhibits choice overload if it violates monotonicity – i.e., the probability of choosing the default option increases as more alternatives are added to the choice set.
-The second is based on the null hypothesis of the random utility model (RUM) – i.e.
-there is a set of utility functions, a (choice set independent) probability distribution over that set, and the probability of choosing any option in any choice set is the probability of the set of utility functions that make that option maximal.
+Item heterogeneity Consider another variant of the IL example: Of the 24 jams IL use, 23 are variants of gooseberry jam, and one is strawberry jam.
+10% of people like gooseberry jam, while everyone likes strawberry jam.
+This means that, absent choice overload, the probability of buying jam from the 24 jam set would be 100%.
+For a smaller choice set, it would be 100% if the strawberry jam were included or 10% otherwise.
+If the smaller set of jams selected by the researched happened not to include strawberry jam, then choice overload would have to be extremely strong to put active choice below the 10% benchmark.
+Example 1.3.
+Tighter bounds Let the choice universe be $X=\{a,b,c,d\}$, where $d$ stands for default.
+Suppose $a,b,c$ are individually chosen across a population with probability $20\%$ from choice sets $\{a,d\},\{b,d\},\{c,d\}$.
+Suppose also that the probability of non-default choice is $40\%$ from any of $\{a,b,d\},\{a,c,d\},\{b,c,d\}$.
+Assuming no indifferences, these probabilities can be reconciled with a Random Utility Model (RUM, precisely defined later); however, they imply that the preferences $a\succ d$, $b\succ d$, and $c\succ d$ occur in disjoint subsets of the population.
+Thus, non-default choice probability from $X$ must be at least $60\%$, even though the largest such probability observed in smaller sets is “only” $40\%$.
+We introduce new tests for choice overload for a data set in which default versus nondefault choice is observed for a grand set $X$ and numerous subsets thereof (always including the same default option).
+We define choice overload as a situation in which the default is chosen too often in $X$ relative to some benchmark.
+We identify two tests based on two different benchmarks.
+The first equates choice overload with violations of monotonicity – that is, the probability of choosing the default option increases as choice sets expand.
+This is the approach previously used in the literature, though the structure of our data still allows for an improved test.
+The second is based on presupposing a RUM – that is, choices can be explained by a (choice set independent) probability distribution over a set of utility functions.
 As example 1.3 illustrates, consistency with RUM implies monotonicity, but not vise versa, so the latter definition provides a more sensitive test for choice overload.
 [^2] In general, it is well known that stochastic monotonicity is necessary but not sufficient for random utility.
-Example 1.2 clarifies that coarsening the domain of observations to active versus passive choice does not fundamentally change this.
-The advantage of the former is that it does not rely on the underlying assumption of random utility - monotonicity is also implied by a number of other models of stochastic choice, such as Manzini and Mariotti’s (2014) Random Consideration Set model.
-If the data set consists only of choice probabilities from $X$ and a single subset $A$ (as is the case in Iyengar and Lepper (2000) and most other studies), then both definitions collapse to the conventional test: Assuming that only the probability of default choice is observed, the data set is consistent with monotonicity and RUM if and only if the probability of choosing $d$ from $X$ is weakly lower than from $A$.
-We therefore propose as a first modification of the standard approach that data should be collected from multiple subsets of $X$.
-In this richer data, monotonicity implies that the choice of the default from $X$ is lower than the lowest probability of choosing $d$ across all observed choice sets $A$.
+Example 1.3 clarifies that coarsening the domain of observations to active versus passive choice does not change this.
+If the data consists only of choice probabilities from $X$ and a single subset $A$, then both definitions coincide and reduce to the current best practice.
+We therefore propose that data should be collected from multiple subsets of $X$.
+In this richer data, monotonicity implies that choice probability of the default from $X$ is less frequent than from any other choice set $A$.
 We call this the Min bound.
-While easy to define, testing the Min bound is subtle because it is a test of multiple hypotheses.
-Even if monotonicity holds, if a finite number of observations are collected from many subsets there is a high probability that one will exhibit lower default choice than is observed from $X$.
-Essentially, the problem is the same as that of the “winners curse” – because the probability of default choice is observed with noise, the minimum of many observations is likely to be below the true minimum.
-To deal with this, we propose tests that take inspiration from closely related problems in the literature on moment inequalities.
-Because monotonicity is necessary but not sufficient for RUM, our model-based definition of choice overload allows us to identify tighter bounds.
-We adapt the approach of McFadden and Richter (1991), who show that random choice data can be rationalized by the RUM if, and only if, a particular linear expression has a solution.
-We refer to the highest default choice probability in the large set for which such a solution exists as the RUM bound.
-The fact that it is choice frequencies, not underlying probabilities, that are observed presents a significant econometric challenge.
-We adapt the nonparametric test of RUM proposed by Kitamura and Stoye (2018), potentially including computational innovations by Smeulders et al.
-(2020) needed for applications with larger sets $X$.
-We apply our tests to a novel experimental data set based on choices from subsets of size $2$ and $3$ of $12$ different choice objects plus a default.
+While easy to define, testing this restriction is subtle because it is a test of multiple hypotheses; we propose an approach that takes inspiration from closely related problems in the literature on moment inequalities.
+We next observe that our model-based definition of choice overload implies tighter bounds.
+We lean on classic work by McFadden and Richter (1991) to characterize the highest default choice probability in the large set that is consistent with the data from subsets and RUM, which we call the RUM bound.
+To implement this condition in practice, we adapt Kitamura and Stoye’s (2018, see also Smeulders et al.
+(2021)) nonparametric test of RUM.
+We apply all tests to a novel experimental data set of choices from subsets of size $2$ and $3$ of $12$ different choice objects plus a default.
 The objects are verbally described sums of numbers as used in Caplin et al.
-(2011); see Figures 1 and 2 for a preview.
-In any given choice set, the subject could choose to stick with a default which provided \$3.50 for sure, or could switch to other available choices.
-Subjects were asked to chose from $10$ such sets, one of which was the default plus the entire set of $12$ other options –henceforth called the grand set– while the other $9$ were uniformly randomly selected from all possible sets with $1$ or $2$ options plus default.
-A total of $2000$ subjects were recruited through Amazon’s Mechanical Turk platform. After removing subjects who failed a comprehension quiz, we were left with $1832$ subjects and $18320$ choices.
+(2011); see Figure LABEL:fig:1 for a preview.
+Subjects were asked to chose from $10$ such sets, one of which was the default plus the entire set of $12$ other options (henceforth called the grand set).
 Traditional approaches would be unlikely to find evidence of choice overload in our data set: The average default choice in small sets is higher than in the grand set, and only 15 of the 78 small sets have a default choice which is significantly lower than that of the grand set.
-Hence if, as in most work in the literature, data was collected from only a single or few small sets, one would be unlikely to detect overload.
-In contrast, even the most ‘brute force’ of our proposed tests –a potentially very conservative but finite-sample valid implementation of the Min bound– do detect it.
-Our test based on RUM provides an unexpected additional insight: While RUM is indeed rejected on the whole data set, it is also rejected if we remove data from choices in the grand set.
-This is because subjects are more likely to choose the default option in choice sets of size $3$ than their data from size $2$ sets and RUM would predict.
-Thus, our results indicate that choice overload type effects can start in choice sets which are much smaller than have previously been demonstrated.
-We hope that our work will have three consequences.
-First, by providing a higher powered test of choice overload, it should clear up the question of whether this is indeed a real phenomenon.
-Second, given that (we suspect) it will show choice overload to be more prevalent than previously thought, we hope it will spur further theoretical and policy work designed to understand its causes and mitigate its effects.
-Finally, by providing a better tool for measuring when choice overload does occur, we hope it will facilitate the above work by providing a better empirical basis on which to theorize.
+In contrast, even the most brute force of our proposed tests – a potentially very conservative but finite-sample valid implementation of the Min bound – do detect it.
+Our test based on RUM provides an unexpected additional insight: While RUM is indeed rejected on the whole data set due to choice overload, it is also rejected if we remove data from choices in the grand set.
+This is because subjects are more likely to choose the default option in choice sets of size $3$ than RUM would predict given their data from size $2$ sets.
+Thus, our results indicate that choice overload type effects can start in choice sets which are much smaller than was previously demonstrated.
+We conclude this section with two important clarifications.
+First, the more sensitive RUM-based test comes at the expense of additional assumptions – for example, monotonicity is also implied by Tversky and Kahneman’s (1991) model of reference dependent preferences.
+So our refined test is really a test of choice overload for a population that is supposed to otherwise adhere to a specific model of rationality.
+We believe, however, that this is in line with what is typically meant by choice overload.
+Secondly, it is important to emphasize that our notion of choice overload is “thin” or behavioral.
+That is, we think of choice overload as a specific feature of choice behavior: that the default is chosen too often in larger sets.
+We do not commit to a theory of what is going on “under the hood.” For example, the increase in default choice set in larger sets observed in our data is plausibly explained by a rational inattention heuristic.
+If true, we would describe this a a case of rational inattention causing choice overload, rather than an alternative description of what is going on.
+We naturally think that “deep” theories of the mechanisms generating choice overload are important and we hope that our work aids future researchers in uncovering them; however, our main aim in this paper is not to substantively advance that inquiry but to provide tools with which to identify choice overload.
 2 Literature Review
-Many studies have examined the phenomenon of choice overload from different perspectives, including different outcome measures and potential moderators.
 Two recent meta-analyses (Chernev et al., 2015; Scheibehenne et al., 2010) report results from 99 experiments in 53 papers and 63 experiments in 50 papers, respectively.
-Given these two reviews, our aim here is not to provide comprehensive coverage of the existing research.
-Instead, we make two points.
-First, to the best of our knowledge, no other paper has made use of the methodology we propose.
-Using the aforecited meta-analyses and Google Scholar, we identified 32 studies from 19 papers that use default choice as a measure of choice overload.
+We refer to these reviews and are content to just make two points.
+First, very few current studies have generated the data needed to perform our tests.
+Using the aforementioned meta-analyses and Google Scholar, we identified 32 studies from 19 papers that use default choice as a measure of choice overload.
 Of these, 20 ask subjects to make choices from a single subset of the grand choice set.
 These studies can do no better than to compare the default choice probability in the small and large choice set.
-The remaining 12 studies ask subjects to make choices from multiple subsets of the large choice set, and so in principle could have implemented either the min bound or tests based on the RUM.
-However, all of them instead compare the average default choice across all small choice sets to that in the larger set, a measure which is necessary but sufficient for the aforementioned Min bound, which in turn is necessary but not sufficient for consistency with RUM.
-Applying our approach to the data from these previous studies (e.g., Chernev (2005), Gingras (2003), or White and Hoffrage (2009)) is an interesting avenue for future research.
-While not explicitly designed to test for choice overload, the experiment of Aguiar et al.
-(2023) does contain the type of data needed to perform our test.
+The remaining 12 studies ask subjects to make choices from multiple subsets of the large choice set, but only collect a small number of choices from each.
+As a result, they instead compare the average default choice across all small choice sets to that in the larger set, a measure which is necessary but not sufficient for the aforementioned Min bound, which in turn is necessary but not sufficient for consistency with RUM.
+While applying our approach to data from these previous studies is an interesting avenue for future research, the small number of observations in each small set make our tests hard to implement.
+[^3] We note that our tests can be applied to any data – not just experimental – where a large sample of default vs non-default choice is observed from a choice set and multiple subsets.
+For example, data on health insurance plan choices, such as that used in Abaluck and Gruber (2023), often has variation in the specific plans offered as well as the number offered which could be used for our analysis.
+We leave such applications of our tests for future research.
+Although not explicitly designed to test for choice overload, the experiments of Aguiar et al.
+(2023) contain the type of data needed to perform our test.
 Choices were observed from all subsets of a grand set of six lotteries, with a default that is always available.
-The authors report finding no evidence of choice overload, and our tests confirm this result.
-[^6] While we performed the test on all their treatments, their ‘low’ complexity treatment is the best fit to apply the test.
-This is because, in the ‘medium’ and ‘high’ treatments, lottery prizes were expressed as sums, and different subjects saw different sums, arguably making them different choice alternatives.
-This may be due to the fact that the default alternative was chosen to be ‘obviously’ worse than the other available alternatives.
+The authors report no evidence of choice overload, and our tests confirm this result.
+This may be due to the fact that the default alternative in their experiment was chosen to be obviously worse than the other available alternatives.
 The second point is that the veracity and scope of choice overload is far from established.
 Some direct replications have failed (Scheibehenne, 2008).
-One meta-analysis (Scheibehenne et al., 2010) finds a mean effect of set size on measures of choice overload to be zero, but noted a high degree of variance.
-A more recent analysis (Chernev et al., 2015) identifies four variables which can increase the incidence of choice overload: decision difficulty (for example due to time constraints), choice set complexity (for example due to hard to compare alternatives), preference uncertainty (for example because the decision maker is unsure how to aggregate their preferences across many dimensions), and decision goal (for example because the decision maker is not really committed to making a purchase).
-Chernev et al.
-(2015) argue that, taking these mediators into account, there is evidence of a robust choice overload effect.
-We believe that the measures we introduce can contribute to this ongoing debate by providing more highly powered evidence on when choice overload is occurring.
-Low powered tests may in part explain some of the contradictory results in the existing literature – for example, some studies may by chance use a small choice set that leads to choice overload, while others do not.
-By making it easier to spot cases of overload, we anticipate movement towards a greater acceptance that choice overload does indeed occur.
+One meta-analysis (Scheibehenne et al., 2010) finds the mean effect of set size on measures of choice overload to be zero, but notes a high variance.
+A more recent analysis (Chernev et al., 2015) identifies four variables which can increase the incidence of choice overload: decision difficulty (for example due to time constraints), choice set complexity (for example due to hard-to-compare alternatives), preference uncertainty (for example because the decision maker is unsure how to aggregate their preferences across many dimensions), and decision goal (for example because the decision maker is not really committed to making a purchase).
+Finally, our testing problem is related to that studied in Kono et al.
+(2023).
+[^4] A recent paper by Turansick (2025) is more distantly related.
+He considers characterizations of RUM when not all choice problems are observable, but for those that are all probabilities are observed.
+We both test RUM when not all choice probabilities from a choice set are observed.
+While also taking inspiration from McFadden and Richter (1991), Kono et al.
+identify a different set of conditions using the Block-Marschak polynomials.
+This approach is elegant and may have computational benefits, but there are some reasons why our approach is more directly applicable to the task at hand.
+First, while Kitamura and Stoye (2018) (and we) use RUMs as motivation, they (and we) really just test whether vectorized choice probabilities lie in a certain polyhedral cone.
+In application to RUM, the vertices of this cone are defined by “choice types” that are conventional utility maximizers, but one can analogously use this approach to test the “random utility extension” of other models of behavior (we do this with generalizations of RUM that capture choice overloaded behavior).
+In addition, Kono et al.
+(2023) assume that the collection of observable choice sets is closed under set expansion; we do not need this and it does not hold in our data.
 3 Theory
-We now present the theoretical underpinnings of our test for choice overload.
-We do so in three stages.
-First, we describe our tests under the assumption that we can perfectly observe the probabilities with which each alternative is chosen in each choice set.
-Second, we discuss the econometric techniques needed to handle the fact that our experimental data are sample frequencies, not population probabilities.
-Finally, we describe the steps necessary to implement these econometric tests, with additional considerations that proved unnecessary for our specific experiment relegated to Appendix B.
+We now present the theoretical underpinnings of our test.
+We initially assume that we can perfectly observe default choice probabilities for each choice set; in a second step, we develop the econometric tools required because our actual data are finite samples.
+Additional computational considerations that proved unnecessary for our experiment are relegated to supplementary materials.
 3.1 The Population-Level Testing Problem
 Let $X$ be a finite set of alternatives and $d$ a default alternative contained in $X$.
-Let $\mathcal{D}\subset 2^{X}/\emptyset$ be a collection of choice sets, all of which contain the default alternative, and which includes the grand set $X$.
-Our data set consists of observations of choice behavior in each set in $\mathcal{D}$ [^7] We follow established terminology in revealed preference theory by calling this a “data set,” but note that statistically speaking, these are population probabilities; inference for finite samples is considered in the next subsection.
-Specifically, we assume that we observe a function $p_{d}:\mathcal{D}\rightarrow[0,1]$, with the interpretation that $p_{d}(A)$ is the probability of choosing the default $d$ from choice set $A$.
-Two notes on our data are in order.
-First, we assume that we observe the population probability with which the default is chosen in each choice set, but not that we can track individuals across choice problems; i.e., we observe a repeated cross section rather than a panel.
-This makes our approach applicable to many between-subject data sets that have this property.
-[^8] We note that the data we collect in our experiment is somewhat richer than this, as it contains multiple observations from the same individual.
-We maintain this assumption to increase the generalizability of our approach.
-Second, we assume that we observe only the probability with which the default was chosen in each choice set, not the probability with which specific non-default options are chosen.
-This is consistent with our desire to to focus on choice overload effects; the limited data only allows us to detect violations of utility maximization that occur to due too much or too little default choice from a set.
-In order to design a test for choice overload, we first need to precisely define the term.
-Here we take two approaches.
-The first is to use an explicitly model-free definition, in the same way that risk aversion can be defined without reference to expected utility.
-An obvious candidate for such a definition is through violations of choice monotonicity, which states that the probability of choosing $d$ cannot increase as more options are added to the choice set.
+Let $\mathcal{D}\subset 2^{X}/\emptyset$ be a collection of choice sets, all of which contain $d$; in most of what follows, and in our experimental data, $\mathcal{D}$ will contain $X$.
+Suppose initially that we observe a function $p_{d}:\mathcal{D}\rightarrow[0,1]$, where $p_{d}(A)$ is the probability of choosing the default $d$ from choice set $A$.
+We assume that we observe the population probability with which the default is chosen in each choice set, but not that we can track individuals across choice problems.
+This makes our approach applicable to many between-subject data sets.
+Crucially, we also assume that we observe only the probability with which the default was chosen in each choice set, not the probability with which specific non-default options are chosen.
+This is consistent with our desire to to focus on choice overload effects; the limited data only allows us to detect violations of utility maximization that occur to due too much or too little default choice from a set, and the tests we will define will identify the former violations as ‘choice overload’.
+We next consider two possible definitions of choice overload. The first one equates it with violations of choice monotonicity.
 Definition 3.1.
-A data set $\{\mathcal{D},p\}$ satisfies monotonicity if, for any $A,B\in\mathcal{D}$ such that $A\subset B$,
+Probabilities $p_{d}$ satisfy monotonicity if, for any $A,B\in\mathcal{D}$ such that $A\subset B$,
 $p_{d}(A)\geq p_{d}(B)$
-Specifically, one could declare choice overload to have occurred if there is a higher probability of default choice in the large choice set than in the smaller one.
-The canonical choice overload experiment, in which $\mathcal{D}=\{A,X\}$, applies this test.
-More generally, one can define
+The canonical choice overload experiment, in which $\mathcal{D}=\{A,X\}$, tests this condition. More generally, one can define
 $p_{d}^{min}(X)=\min_{A\in\mathcal{D}\setminus X}p_{d}(A)$
 as the smallest observed probability of choosing $d$ in any set other than $X$.
-If $p_{d}(X)>p_{d}^{min}(X)$ then monotonicity is violated and the data set exhibits choice overload.
-We refer to $p_{d}^{min}(X)$ as the Min bound.
-Data that violates the Min bound is inconsistent with a number of models – most obviously RUM, but also the stochastic consideration set model of Manzini and Mariotti (2014) [^9] Other, more general models of stochastic consideration do allow for choice overload type effects - see for example Cattaneo et al.
-(2020, 2021) and models of reference dependent preferences such as Tversky and Kahneman (1991).
+Monotonicity is violated iff $p_{d}(X)>p_{d}^{min}(X)$.
+We therefore refer to $p_{d}^{min}(X)$ as the Min bound.
+We will say that data that violates this condition as exhibiting choice overload with respect to the Min bound.
+Such data is inconsistent with a number of models – most obviously RUM, but also models of reference dependent preferences such as Tversky and Kahneman (1991).
+[^5] This is also true for the stochastic consideration set model of Manzini and Mariotti (2014), a special case of RUM.
+Other more general models of stochastic consideration allow for choice overload type effects – see, for example, Cattaneo et al.
+(2020, 2021).
 A second approach is to define choice overload as a violation of a specific model.
-Here, the most obvious candidate would be RUM, and we will work with it, although the basic idea would generalize to any model that we know how to test.
+Here, the most obvious candidate is RUM, and we will work with it, although the basic idea would generalize to any model that we know how to test.
 Thus, call a data set stochastically rationalizable if it could have been generated by a RUM.
 Then we can think of it as revealing choice overload if it would be stochastically rationalizable except that the probability of default choice in larger choice sets is too large.
+The basic idea is that one would declare a data set to exhibit choice overload if the default is chosen too often in the grand set $X$, given the choice probabilities from the smaller sets and the constraints imposed by RUM.
+One could implement this either by testing for the validity of RUM both including or excluding data from $X$, or by computing counterfactual bounds on $p_{d}(X)$ as the set of all default choice probabilities on $X$ that would be rationalizable jointly with the other observed probabilities.
 To illustrate this idea, consider the following definitions.
-Note that, while $\mathcal{D}$ in our application contains $X$, these definitions refer to a generic $\mathcal{D}$ that may or may not do so.
 Definition 3.2.
-A data set $\{\mathcal{D},p\}$ is consistent with RUM if there exist a finite collection $\mathcal{U}$ of one-to-one utility functions on $X$ and a probability distribution $\rho\in\Delta(\mathcal{U})$ such that, for every $A\in\mathcal{D}$,
+Probabilities $p_{d}$ are consistent with RUM if there exist a finite collection $\mathcal{U}$ of one-to-one utility functions on $X$ and a probability distribution $\rho\in\Delta(\mathcal{U})$ such that, for every $A\in\mathcal{D}$,
 $p_{d}(A)=\sum_{u\in\mathcal{U}\mid d=\arg\max u(A)}\rho(u)$
-For any set $A\subseteq X$, we can then define a maximal bound on the choice of default using the default choice probabilities from the subsets of $A$ and consistency with RUM.
-The basic idea here goes back to Varian (1982, 1983): A counterfactual choice behavior is in the predictive bounds if, and only if, that choice behavior and previously observed ones are jointly rationalizable.
-Formally, we have:
+For any $A\subseteq X$, we can then define a maximal bound on the choice of default using the default choice probabilities from the subsets of $A$ and consistency with RUM.
+The basic idea here goes back to Varian (1982, 1983): A counterfactual choice behavior is in the predictive bounds if, and only if, that choice behavior and previously observed ones (in our case, behavior on small choice sets) are jointly rationalizable.
+Formally:
 Definition 3.3.
-For a dataset $\{\mathcal{D},p\}$ and choice problem $A\subseteq X$, define
+Define
 $p_{d}^{RUM}(A)=\sup x\in[0,1]$
-such that the data set
+subject to probabilities
 $\tilde{p}_{d}(\tilde{A})=\left\{\begin{array}[]{c}x\text{ if }\tilde{A}=A\\ p_{d}(\tilde{A})\text{ otherwise}\end{array}\right.$
-defined on the choice sets $\mathcal{D}_{A}\equiv\{\tilde{A}\in\mathcal{D}:\tilde{A}\subset A\}$ is consistent with RUM.
-For most of this paper, we say that a data set exhibits choice overload if $p_{d}(X)>p^{RUM}_{d}(X)$.
+on choice sets $\mathcal{D}_{A}\equiv\{\tilde{A}\in\mathcal{D}:\tilde{A}\subseteq A\}$ being jointly consistent with RUM.
+For most of this paper, we say that a data set exhibits choice overload according to the RUM bound if $p_{d}(X)>p^{RUM}_{d}(X)$.
 However, the definition of $p_{d}^{RUM}(\cdot)$ allows for choice overload to “kick in” for smaller choice sets and we will consider that later.
-Conceptually, the Min bound and the RUM bound deal with different reasons why observations from a single small choice set might not effectively identify choice overload.
-The first is heterogeneity in the quality of available items.
-Consider a grand choice set that consists of a number of not very appealing jam flavors (gooseberry, bark, salmon) and one extremely appealing flavor (strawberry).
-Absent any choice overload, we would expect to see high levels of default choice in small sets that do not include the strawberry jam, and low levels of default choice in both small sets that included the strawberry jam, and the grand set.
+The Min and RUM bounds speak to different reasons why observations from a single small choice set might not effectively identify choice overload.
+The first is heterogeneity in the quality of available items, as illustrated in example 1.2 .
+Consider a grand choice set that consists of a number of not very appealing jam flavors and one extremely appealing flavor (strawberry, say).
+Absent any choice overload, we would expect high levels of default choice in small sets that do not include the strawberry jam, and low levels of default choice in both small sets that included the strawberry jam, and the grand set.
 Thus, if a researcher randomly selected for analysis a small choice set without the strawberry jam, it would make it very hard to spot choice overload.
-Collecting data on all small choice sets and applying the Min bound would help to solve this problem, because the researcher would observe choice from a small set in which strawberry jam was included.
-A second issue is preference heterogeneity.
-Consider example 1.2 from the introduction.
-In this case, all alternatives are ex ante identical, so there is no heterogeneity in item quality.
-This in turn means that the default choice probability is the same in any choice set with the same cardinality.
-Thus, collecting (for example) data from multiple sets of size two and applying the Min bound does not help.
-However, as the example demonstrates, the RUM bound based on observations from choice sets of size two and three does put further constraints on the data.
-This is because the assumption that randomness is driven by preference heterogeneity puts extra structure on the allowed choice probabilities.
-Thus the RUM-based definition of choice overload gives potentially tighter bounds with which to identify choice overload.
-[^10] Note that, without preference heterogeneity, the RUM bound will equal the Min bounds.
-If all subjects have the same preferences, then in all small choice sets the default choice probability must be either zero or one.
-Thus, the Min bound must either be zero, if in some choice set the default is never chosen, or one, if the default is always chosen in every choice set.
-In both cases the RUM bound would be the same as the Min bound.
-A disadvantage of RUM bound is that it requires data on $\mathcal{D}/X$ to be consistent with RUM.
+Collecting data on all small choice sets and applying the Min bound would address this problem.
+The second issue is preference heterogeneity.
+Consider Example 1.3: Here, all alternatives are equally likely to be attractive.
+Default choice probability is the same in all choice sets of the same size, and so a simple application of the Min bound will not increase power to detect choice overload.
+However, the RUM bound based on observations from choice sets of size two and three further constrains choice probabilities.
+(In a perfectly homogeneous population, the bounds coincide.)
+A feature of the RUM bound is that it requires choice from $\mathcal{D}/X$ to be consistent with RUM.
 There are two possible issues with this.
-First, it could be that the population choice distributions on on $\mathcal{D}/X$ are inconsistent with RUM; in that case, $p_{D}^{RUM}(X)$ is not well-defined.
-Second, it could be that rationalizable population distributions generate non-rationalizable finite sample frequencies; in that case, one could define a feasible version of $p_{d}^{RUM}(X)$, and we will do so in Section 4.3.
+First, it could fail at the population level; in that case, $p_{D}^{RUM}(X)$ is not well-defined.
+Second, rationalizable population distributions will still, at least occasionally, generate non-rationalizable finite sample frequencies; in that case, one could define a feasible version of $p_{d}^{RUM}(X)$, and we will do so in Section 4.2.
+Notice that this feature of the RUM bound, along with our focus only on default choice probabilities, means that many well known behavioral phenomena will not show up as choice overload as per our definition – either because they would lead to violations of RUM in the set $\mathcal{D}/X$, or because they would not be observable due to the coarsening of our data.
+For example, consider a case in which the DM was first asked to choose between two hard to compare alternatives – $x$ and $y$, say – plus a default, and then asked to choose from the same set with an alternative $z$ that is dominated by $x$ but not $y$.
+The asymmetric dominance effect might push people to switch their choices from $y$ to $x$ in the larger choice set, but if this does not affect the choice of default, it will not show up as choice overload.
 It remains to clarify how we can test stochastic rationalizability of data.
-In its essence, this question has been resolved in McFadden and Richter (1991).
-Because we adapt their approach to an environment in which rationalizability is not characterized by standard revealed preference axioms (although they mention the possibility of this generalization) and to keep the treatment self-contained, we will briefly explain the answer.
-The basic insight is that probabilistic choice data can be rationalized if, and only if, it can be expressed as convex combination of data that would be produced by deterministic utility maximizers.
-[^11] As is often the case with seminal results, with ample hindsight the intuition informs a much easier than the original proof (Stoye, 2019).
+For the case of standard stochastic choice data, this question has been resolved by McFadden and Richter (1991, see () for a short proof).
+Here we adapt their approach to our data set.
+The basic insight is that choice probabilities can be rationalized if, and only if, they can be expressed as convex combination of data that would be produced by deterministic utility maximizers.
 To make this precise, construct a matrix $\bm{A}$ s.t.
-each row of $\bm{A}$ corresponds to a given alternative within a particular choice set (i.e., each alternative appears once for every choice set containing it), and each rationalizable deterministic choice pattern corresponds to exactly one column of $\bm{A}$.
-For example, consider a data set consisting of choices from $\left\{a_{1},d\right\}$, $\left\{a_{1},a_{2},d\right\}$ and $\left\{a_{1},a_{2},a_{3},d\right\}$.
-The first row of $\bm{A}$ indicates the choice of $a_{1}$ from $\{a_{1},d\}$, the second row the choice of $d$ from $\{a_{1},d\}$ and so on.
-One column of $\bm{A}$ (namely, the first one in 3.1 below) then represents the choices of someone who picked $a_{1}$ from all three choice sets, which can be rationalized as the choices of a decision maker who prefers $a_{1}$ over all other alternatives in $X$.
-The remaining columns of $\bm{A}$ represent all other rationalizable choice patterns.
-Some book-keeping reveals that in this example, one has: [^12] The size of $\bm{A}$ increases extremely rapidly with the complexity of data; see Kitamura and Stoye (2018) for numerical examples.
-$\begin{array}[]{c}a_{1}|\left\{a_{1},d\right\}\\ d|\left\{a_{1},d\right\}\\ a_{1}|\left\{a_{1},a_{2},d\right\}\\ a_{2}|\left\{a_{1},a_{2},d\right\}\\ d|\left\{a_{1},a_{2},d\right\}\\ a_{1}|\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{2}|\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{3}|\left\{a_{1},a_{2},a_{3},d\right\}\\ d|\left\{a_{1},a_{2},a_{3},d\right\}\end{array}\left\{\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&0&0&0&1&1&1&1\\ 1&1&0&0&0&0&0&0\\ 0&0&1&1&0&0&1&1\\ 0&0&0&0&1&1&0&0\\ 1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&1&0\\ 0&1&0&1&0&1&0&1\\ 0&0&0&0&1&0&0&0\end{array}\right\}=\bm{A}.$ (3.1)
-The core insight of McFadden and Richter (1991) is summarized in the following theorem.
+each row of $\bm{A}$ corresponds to a given alternative within a particular choice set (i.e., each alternative appears once for every choice set containing it).
+Each column corresponds to a deterministic choice pattern rationalizable by a different (strict) preference profile over $X$.
+For example, let $\mathcal{D}=\{\{a_{1},d\},\{a_{1},a_{2},d\},\{a_{1},a_{2},a_{3},d\}\}$ and let the first row of $\bm{A}$ indicate choice of $a_{1}$ from $\{a_{1},d\}$, the second row choice of $d$ from $\{a_{1},d\}$, and so on.
+One column of $\bm{A}$ (namely, the first one in (3.1)) then represents the choices of someone who picked $a_{1}$ from all three choice sets, rationalizable by preferences ranking $a_{1}$ first.
+Constructing the remaining columns from all other rationalizable choice patterns, one has:
+$\begin{array}[]{c}a_{1}\mid\left\{a_{1},d\right\}\\ d\mid\left\{a_{1},d\right\}\\ a_{1}\mid\left\{a_{1},a_{2},d\right\}\\ a_{2}\mid\left\{a_{1},a_{2},d\right\}\\ d\mid\left\{a_{1},a_{2},d\right\}\\ a_{1}\mid\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{2}\mid\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{3}\mid\left\{a_{1},a_{2},a_{3},d\right\}\\ d\mid\left\{a_{1},a_{2},a_{3},d\right\}\end{array}\left\{\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&0&0&0&1&1&1&1\\ 1&1&0&0&0&0&0&0\\ 0&0&1&1&0&0&1&1\\ 0&0&0&0&1&1&0&0\\ 1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&1&0\\ 0&1&0&1&0&1&0&1\\ 0&0&0&0&1&0&0&0\end{array}\right\}=\bm{A}.$ (3.1)
+McFadden and Richter’s (1991) core insight is the following.
 Theorem 3.1.
-Let $\pi$ be the observed choice probabilities associated with each row of the matrix $\bm{A}$ derived from some stochastic choice data set.
-That data set is rationalizable by RUM if and only if there exists a vector $\nu\in\Delta^{H-1}$ (the $(H-1)$-dimensional unit simplex, where $H$ is the number of columns of $\bm{A}$) such that
-$\bm{A}\nu=\pi$
-In order to adapt this approach to our idealized data set, we need to deal with the fact that we only observe whether or not the default was chosen.
-To do so, we premultiply the matrix $\bm{A}$ with a matrix $\bm{B}$ that merges events we do not separately observe – i.e.
-the choice of different non-default options.
-In the above example, we would have
-$\bm{B}=\left\{\begin{array}[]{ccccccccc}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&1&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right\}$
-A corollary to Theorem 3.1 then characterizes $p_{d}^{RUM}(X)$:
+Let the vector $\rho$ collect observed choice probabilities in order corresponding to the rows of $\bm{A}$.
+These probabilities are rationalizable by RUM if, and only if, there exists a vector $\nu\in\Delta^{H-1}$ (the $(H-1)$-dimensional unit simplex, where $H$ is the number of columns of $\bm{A}$) such that
+$\bm{A}\nu=\rho.$
+We slightly adapt this approach because we assume that we only observe whether $d$ was chosen from any choice set.
+Thus, premultiply $\bm{A}$ by a matrix $\bm{B}$ that merges the choice of different non-default options in a given menu.
+Columns of $\bm{B}\bm{A}$ then represent different rationalizable deterministic default vs non-default choice patterns.
+Each choice set in $\mathcal{D}$ is represented by two rows in $\bm{B}\bm{A}$, the first of which indicates non-default choice from the choice set, and the second representing default choice.
+For the example above, $\bm{B}$ and $\bm{BA}$ are as follows:
+$\bm{B}=\left\{\begin{array}[]{ccccccccc}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&1&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right\},\quad\bm{BA}=\left\{\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&0&0&0&1&1&1&1\\ 1&1&1&1&0&0&1&1\\ 0&0&0&0&1&1&0&0\\ 1&1&1&1&0&1&1&1\\ 0&0&0&0&1&0&0&0\end{array}\right\}$
+A corollary to Theorem 3.1 then characterizes $p_{d}^{RUM}(X)$.
+To state it formally, let the vector $\pi$ collect probabilities of active and passive choice in order corresponding to rows of $\bm{BA}$; in our example,
+$\pi=\begin{pmatrix}1-p_{d}(\{a_{1},d\})\\ p_{d}(\{a_{1},d\})\\ 1-p_{d}(\{a_{1},a_{2},d\})\\ p_{d}(\{a_{1},a_{2},d\})\\ 1-p_{d}(\{a_{1},a_{2},a_{3},d\})\\ p_{d}(\{a_{1},a_{2},a_{3},d\})\end{pmatrix}.$
+Then we can write:
 Corollary 3.1.
-Let $\pi$ be the observed choice probabilities associated with each row of the matrix $\bm{BA}$ derived from a data set $\{\mathcal{D},p\}$ which reports only the probability of default choice.
-That data set is rationalizable by RUM if and only if their exists a probability vector $\nu$ such that
-$\bm{BA}\nu=\pi.$
+A probability vector $\pi$ as just defined is rationalizable by RUM if, and only if, there exists $\nu\in\Delta^{H-1}$ such that
+$\bm{BA}\nu=\pi.$ (3.2)
 Further, let $\bm{a}$ be the row of $\bm{A}$ that corresponds to default choice from $X$, then
@@
 For intuition, observe that in (3.1), the vector $\bm{a}$ is the last row of $\bm{A}$.
-Equivalently, it is the indicator of the choice type being the one whose choice is always the default (this is the only type who chooses the default from $X$).
-The bound simply maximizes the probability corresponding to this type, subject to overall data being stochastically rationalizable.
-In practice, $p_{d}^{RUM}(X)$ as stated is well-defined only if $\pi$ is stochastically rationalizable.
-Since the set of rationalizable probability vectors is “small” (we will elaborate on this in Section 4.3), empirical choice frequencies will typically –and do in our data– fail this.
+Equivalently, it is the indicator of the unique choice type whose choice is always the default.
+The bound simply maximizes the probability of this type, subject to overall data being stochastically rationalizable.
+[^6] Expression (3.3) presupposes that $\pi$ is stochastically rationalizable.
+Since the set of rationalizable probability vectors is “small” (we will elaborate on this in Section 4.2), empirical choice frequencies may fail this.
 In that case, a feasible version of the bound can be computed by substituting a constrained estimator of $\pi$.
 This will be illustrated later.
-We finally note that the above machinery opens the door to using this testing approach on other, for instance non-RUM, models.
-Nothing in the above development forces $\bm{A}$ to contain columns that correspond exactly to conventionally rationalizable behaviors.
+Note that, when we only observe default vs non-default choice, multiple deterministic rational choice types may be indistinguishable – e.g., the first four columns of $\bm{A}$ represent types that never choose the default from choice problems in $\mathcal{D}$ and hence are identical in $\bm{BA}$.
+In practice, one may simplify problem (3.2) by eliminating such repetitions.
+We finally note that we can apply this testing approach to other (e.g.
+non-RUM) models.
+Nothing forces $\bm{A}$ to contain columns that correspond exactly to conventionally rationalizable behaviors.
 By adding (removing) columns of $\bm{A}$, one can test less (more) restrictive models.
-Indeed, in our empirical application we will use this approach to test:
+We will use this observation as the basis of method for identifying choice overload. We will empirically test:
   • (i)
@@
   • (ii)
-    A model which admits all the rationalizable deterministic choice patterns of RUM but also allows for choosing $d$ from choice set $X$ (regardless of behavior on smaller sets).
+    A relaxation of RUM that allows for choosing $d$ from choice set $X$, regardless of behavior on smaller sets.
   • (iii)
-    The same model as (ii), except that a choice type may also switch to choosing $d$ for all choice sets of cardinality $3$.
-(Types that choose $d$ from all sets of cardinality $3$ must also choose $d$ from choice set $X$.)
-Models (ii) and (iii) capture ‘choice overloaded’ behavior in that the only violations of standard rationality they allow are switches to the default when the choice set expands.
-Moving from (i) to (iii) enlarges $\bm{A}$, and therefore increases computational cost, but does not add conceptual difficulties.
-Since the resulting models are nested (listed in increasing order of generality), one can then ask what is the least permissive model that is not rejected in the data.
-This will allow us to unpack what sort of choice overloaded behavior, if any, could have generated our data; for reasons that will become clear, this line of analysis will be central to our results.
-As a preview, model (iii) will be the only one not rejected in our data.
-This analysis is in Section 4.3, where we also argue that model (iii) is itself a restrictive model.
-We note that a further generalization which captures choice overload would be to allow switches to the default not as a function of set cardinality (as in (ii) and (iii)), but along any distinct path of choice set expansion.
-Dean (2008) considers a similar model.
-We do not consider this generalization due to computational complexity (the corresponding $\bm{A}$-matrix is very large) and because we know it would not be rejected in our data given it nests (iii).
+    The same model as (ii), except that a choice may switch to $d$ for all choice sets of cardinality $3$. (Types that choose $d$ from all sets of cardinality $3$ must also choose $d$ from $X$.)
+The RUM bound is violated if (i) is rejected while model (ii) is not: Model (ii) captures choice overloaded behavior by only allowing violations of standard rationality through switching to $d$ when the choice set expands.
+Model (iii) expands on the standard notion of choice overload by allowing it to “kick in” at smaller choice sets.
+Moving from (i) to (iii) enlarges $\bm{A}$ but does not add conceptual difficulties.
+Since the resulting models are nested, one can then ask what is the least permissive model that is not rejected in the data.
+This will allow us to unpack what sort of choice overloaded behavior, if any, could have generated our data.
+In Section 4.2, we show that model (iii) is the only one not rejected in our data and argue that is still a restrictive model.
 3.2 Econometric Tests
-We next explain some testing strategies that connect the above ideas to recent advances in econometrics.
-To this purpose, we consider samples that were generated by randomly drawing individuals and then exposing each individual to an (i.i.d.
-randomly generated) selection of choice problems, i.e.
-subsets of $\mathcal{D}$ [^13] This mirrors our empirical design, which was partly chosen because, unlike stratified sampling or mean-reverting coins, it is easy to bootstrap.
+We next explain testing strategies that connect the above ideas to recent advances in econometrics.
+[^7] This section can be skipped without loss of continuity.
+To this purpose, we consider samples that were generated by randomly drawing individuals and then giving each individual a (i.i.d.
+randomly generated) selection of choice problems.
+[^8] This mirrors our empirical design, which was partly chosen because, unlike stratified sampling or mean-reverting coins, it is easy to bootstrap.
 In particular, data may contain choices from the same individuals in different choice sets, as is the case in our application.
-Note also that in this section, the term “data” is used in the statistical sense, i.e.
-referring to a finite sample from an underlying population.
-We begin with two tests of the Min bound. We will estimate choice probabilities $p_{d}(\cdot)$ by the analogous sample frequencies, which can be informally defined as
-$\hat{p}_{d}(A)\equiv\frac{\sum\bm{1}\{\text{choice problem is }A,\text{ choicemade is }d\}}{\sum\bm{1}\{\text{choice problem is }A\}},$ (3.3)
-where $\bm{1}\{\cdot\}$ is the indicator function and the sums are taken over all data points.
-For all but the finite-sample test that we present first, any estimator of $p_{d}(\cdot)$ whose asymptotic distribution is normal or approximated by the simple nonparametric bootstrap would suffice.
-Sample averages have both properties as long as they are not close to degenerate; by direct inspection, this condition is easily met in our data.
+We estimate choice probabilities $p_{d}(\cdot)$ by the analogous sample frequencies $\hat{p}_{d}(A)$.
+For all but the finite sample test that we present first, any estimator of $p_{d}(\cdot)$ whose asymptotic distribution is normal or approximated by the simple nonparametric bootstrap would suffice.
+Sample frequencies have both properties as long as they are not close to degenerate, where “close to” is relative to sample size.
+We chose a relatively large sample size precisely to ensure this, and it easily holds in our data.
 3.2.1 A Finite-Sample Test of the Min Bound
@@
 testing whether $p_{d}(X)\leq p_{d}(A)$ for a specific $A\subset X$.
-To this purpose, define $\hat{p}_{d,A}(X)$ in exact analogy to (3.3) but dropping observations from subjects who also saw choice problem $A$.
-In words, we estimate $p_{d}(X)$ by its simple empirical analog among subjects who did not see $A$.
+To this purpose, define the “leave-$A$-out” sample frequency $\hat{p}_{d,A}(X)$ by dropping observations from subjects who also saw choice problem $A$.
 As a result, $\hat{p}_{d}(A)$ and $\hat{p}_{d,A}(X)$ estimate binomial proportions in two mutually exclusive samples.
-We will therefore apply Fisher’s (1992) exact test for binomial proportions to $H_{0}:p_{d}(X)\leq p_{d}(A)$ for any given $A$.
-Of course, we need to account for the fact that we will conduct many such tests once ($78$ in our application) and cannot assume independence.
+We therefore apply Fisher’s (1992) exact test for binomial proportions to $H_{0}:p_{d}(X)\leq p_{d}(A)$ for any given $A$.
+Of course, we need to account for the fact that we conduct many such tests once ($78$ in our application) and cannot assume independence.
 Our first approach is Bonferroni adjustment, that is, all p-values are multiplied by $78$.
 An advantage of this approach is that it ensures finite-sample (as opposed to asymptotic) size control.
-However, its power is limited through three channels: The estimators $\hat{p}_{d,A}(\cdot)$ discard data; Fisher’s exact test is in general conservative due to integer issues; and Bonferroni adjustment is conservative.
+However, its power is limited through three channels: The estimators $\hat{p}_{d,A}(\cdot)$ discard data; Fisher’s exact test is in general conservative due to integer issues and a strong sense of conditional validity; Bonferroni adjustment is conservative.
 In practice, with the sample sizes that we generated for our empirical application, we expect only the last channel to have an appreciable effect.
@@
 A common concern with such adjustments is that, if the results of some of these tests appear obvious, one might needlessly lose power.
-Indeed, consider the visualization of our data in Figure 3: The default probabilities in the right-hand cluster are obviously much higher than in the grand choice set (corresponding to the dashed line).
+Indeed, in our experimental data, the default probabilities in a number of choice sets are obviously much higher than in the grand choice set.
+[^9] Figure 7 in Supplemental Appendix A.2 provides a visualization of this.
 Can we restrict attention to only those inequality conditions that might reasonably bind?
-This question has received considerable attention in the econometric literature on moment inequalities and is by now well understood.
+This question has received considerable attention in the econometric literature on moment inequalities.
 We implement a method that can be seen as special case of Andrews and Soares (2010) and also of Chernozhukov et al.
 (2013), both of whom establish its validity under rather general conditions.
-The method can in principle use many test statistics; for concreteness, set
+The method can use many test statistics; for concreteness, set
 $t=\hat{p}_{d}(X)-\min_{A\in\mathcal{D}\setminus X}\hat{p}_{d}(A).$
@@
 This parameter cannot be pre-estimated with sufficient accuracy [^10] Technically, it enters the asymptotic distribution scaled by $\sqrt{n}$.
-Our exposition is slightly informal because the issue is well understood in the literature; see also Canay and Shaikh (2017, section 4) for a survey.
+See Canay and Shaikh (2017, section 4) for a survey.
 and so we must conservatively approximate it.
@@
 we (i.i.d.
-uniformly with replacement) resample individuals and use all responses from a given resampled individual; this ensures that correlation patterns in $(\hat{\pi}-\pi)$ due to eliciting several responses per individual are captured.
+uniformly with replacement) resample individuals and use all responses from a given resampled individual; this ensures that correlation patterns in $\hat{p}_{d}(\cdot)$ due to eliciting several responses per individual are captured.
   2. 2.
     Use a pre-test with size converging to 00, e.g.
-$\alpha_{n}=1/\log(n)$, and discard from consideration any sets $A$ s.t.
+$\alpha_{n}=\alpha/\log(n)$, where $\alpha$ is the test’s nominal size.
+Discard from consideration any sets $A$ s.t.
 the null hypothesis $H_{0}:p_{d}(X)\geq p_{d}(A)$ is rejected at significance level $\alpha_{n}$.
@@
 Second, the test may be extremely conservative if we accordingly recenter all $78$ estimators.
-Therefore, we pre-screen choice items whose default probability is likely to much exceed $p_{d}(X)$.
-This is a special case of Generalized Moment Selection (Andrews and Soares, 2010).
-In particular, the size of the pre-test must go to 00 in order for us to claim that we will asymptotically select all binding constraints.
+Therefore, we pre-screen choice items whose default probability is likely to much exceed $p_{d}(X)$.In particular, because the size of the pre-test goes to 00, we will asymptotically select all binding constraints.
 [^11] Our depiction of the method is simplified by picking a specific test statistic and also filling in specific values for several tuning parameters.
-For example, the sum $\sum_{A\in\mathcal{D}^{*}\setminus X}\max\{\hat{p}_{d}(X)-\hat{p}_{d}(A),0\}$ of constraint violations might yield a test statistic that is more powerful against small but uniform violations.
 Also, rather than letting the size of a pre-test vanish, one could use Bonferroni correction to “spend” some “coverage budget” on the pre-test (Andrews and Barwick, 2012; Romano et al., 2014).
-Again, we omit these generalizations for conciseness because they are known in the literature.
-The test that we implement is as stated, and the rejection in our empirical application is so resounding that none of these choices would plausibly have impacted it.
+This would make no difference in our empirical application because its p-values are far from conventional thresholds.
 3.2.3 An Asymptotic Test of RUM and its Generalizations
@@
 (2021).
-It has seen application to observational data (Deb et al., 2022) as well as lab experiments (Aguiar et al., 2022).
-Because we will slightly adapt the approach to our setting, and also for self-contained exposition, we next briefly explain it.
-We also remind the reader that RUM is here meant in a somewhat loose sense: It is not essential that the population is modelled as mixing a finite number of “admissible” types encoded in the columns of $\bm{A}$, not that admissibility coincides with standard economic rationality; hence, we can use the machinery to test nonstandard and, in particular, nested models.
+It has seen application to observational data (Deb et al., 2023) as well as lab experiments (Aguiar et al., 2023).
+Importantly, the approach only requires that the population is modeled as mixing a finite number of “admissible” types encoded in the columns of $\bm{A}$, not that admissibility coincides with standard economic rationality.
+Hence, we can use the machinery to test nonstandard and, in particular, nested models.
 The Hypothesis Test
-The null hypothesis
+A main insight in Kitamura and Stoye (2018) is that the null hypothesis
 $H_{0}:\bm{BA}\nu=\pi,\exists\nu\in\Delta^{H-1}$
 can equivalently be written as
-$H_{0}:$ $\bm{BA}\nu=\pi,\exists\nu\geq 0$
-$\Longleftrightarrow$ $\pi\in\mathcal{C}\equiv\{\bm{BA}\nu:\nu\geq 0\}$
-$\Longleftrightarrow$ $\min_{\nu\geq 0}\{(\pi-\bm{BA}\nu)^{\prime}\Omega(\pi-\bm{BA}\nu)\}=0,$
+$H_{0}:\min_{\nu\geq 0}\{(\pi-\bm{BA}\nu)^{\prime}\Omega(\pi-\bm{BA}\nu)\}=0,$
 where $\Omega$ is an arbitrary positive definite (and in practice diagonal) weighting matrix.
-Here, the first step is true because a vector $\nu$ can fulfil $\bm{BA}\nu=\pi$ only if its elements sum to $1$, thus we can relax the constraint on $\nu$.
-The last expression states that the residuals from projecting $\pi$ onto the cone $\mathcal{C}$ of rationalizable probabilities must equal 00 and suggests the scaled norm of the corresponding sample residuals as natural test statistic.
+That is, the residuals from projecting $\pi$ onto the cone $\mathcal{C}$ of rationalizable probabilities must equal 00.
+This suggests the scaled norm of the corresponding sample residuals as test statistic.
 Noting the similarity to specification tests in multiple equation models (Sargan, 1958; Hansen, 1982), call this statistic
 $J_{n}\equiv n\min_{\nu\geq 0}\{(\hat{\pi}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}-\bm{BA}\nu)\},$
-where $\hat{\pi}$ stacks estimated choice probabilities $(\hat{p}_{d}(A),1-\hat{p}_{d}(A))$ in an order corresponding to $\pi$ and $n$ is sample size.
-The asymptotic distribution of $J_{n}$ is delicate to estimate because it depends discontinuously on where on $\mathcal{C}$ the true $\pi$ is.
-Valid inference therefore relies on a bootstrap procedure that not only recenters the empirical distribution to be on $H_{0}$ (as in the previous test) but that also shrinks $H_{0}$ in a locally unfavorable direction.
-Heuristically, the idea is once again similar to Generalized Moment Selection in moment inequalities, but the details differ because, while $\mathcal{C}$ is a finite polytope, we can only characterize it in terms of its vertices (the columns of $\bm{A}$) as opposed to its faces (which would correspond to linear inequality constraints and allow for Generalized Moment Selection if we had them).
-Operationally, we recenter the bootstrap distribution of data at the projection onto the tightened cone
-$\mathcal{C}_{\tau_{N}}\equiv\{\bm{BA}\nu:\nu\geq\bm{1}\cdot\tau_{n}/H\},$
-where $\tau_{n}$ is a tuning parameter that we set equal to $\sqrt{\log(\underline{n})/\underline{n}}$ with $\underline{n}$ being expected sample cell size for any but the universal choice problem; $\bm{1}$ is a vector of $1$’s.
-[^16] Simple algebra reveals that $\underline{n}=\tfrac{2qn}{k(k+1)}$, where $k$ is the number of nondefault items in $X$ and $q$ is the number of choice problems other than the grand problem faced by each subject.
-Recall also that $H$ is the length of $\nu$, thus division by $H$ ensures that the above constraint is scaled by $\tau_{n}$ and not by the testing problem’s complexity.
-The essential feature here is that $\tau_{n}$ vanishes but slowly compared to estimation uncertainty in $\hat{\pi}$.
-Then the bootstrap analog of $J_{n}$ is defined as
+where $\hat{\pi}$ is the sample analog of $\pi$ and $n$ is sample size.
+Despite the superficial similarity to well-established methods, the asymptotic distribution of $J_{n}$ is hard to estimate because it depends discontinuously on where on $\mathcal{C}$ the true $\pi$ is.
+However, one main contribution of Kitamura and Stoye (2018) is precisely to overcome this problem.
+Following them, we approximate the distribution of $J_{n}$ by the one of a modified bootstrap analog
 $J^{*}_{n}$$\equiv$ $\min_{\nu\geq\bm{1}\cdot\tau_{n}/H}\{(\hat{\pi}_{\tau_{n}}^{*}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}_{\tau_{n}}^{*}-\bm{BA}\nu)\}$ (3.4)
 $\hat{\pi}_{\tau_{n}}^{*}$$\equiv$ $\hat{\pi}^{*}+\hat{\eta}_{\tau_{n}}-\hat{\pi}$
-$\hat{\eta}_{\tau_{n}}$$\equiv$ $\arg\min_{\nu\geq\bm{1}\cdot\tau_{n}/H}\{(\hat{\pi}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}-\bm{BA}\nu)\}$ (3.5)
-and $\hat{\pi}^{*}$ is a simple nonparametric (clustered, as explained earlier) bootstrap analog of $\hat{\pi}$.
-Validity of the resulting test under reasonable assumptions is established in Kitamura and Stoye (2018).
-We go beyond a completely straightforward implementation of their test because we do not weight questions equally.
-This possibility is anticipated in Kitamura and Stoye (2018) because they require only a diagonal weighting matrix $\Omega$ but has not, to our knowledge, been implemented before.
-It is motivated by the fact that, in our data, choice probabilities pertaining to the universal choice set $X$ will be estimated from a much larger sample cell, and therefore much more precisely, than others.
-Specifically, both the expected and also realized average relative cell size of the last versus other questions are determined by the experimental design, and it is this number that we will use, i.e.
-the weighting matrix will be
-$\Omega=\left(\begin{array}[]{rrrrr}1&\cdots&0&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ 0&\cdots&1&0&0\\ 0&\cdots&0&w&0\\ 0&\cdots&0&0&w\end{array}\right),$
-where $w$ is easy to compute and takes a value of $w\approx 9$ in our experimental design.
-[^17] Indeed, $w=\tfrac{k(k+1)}{2q}$, with $(k,q)$ as in the preceding footnote.
+$\hat{\eta}_{\tau_{n}}$$\equiv$ $\arg\min_{\nu\geq\bm{1}\cdot\tau_{n}/H}\{(\hat{\pi}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}-\bm{BA}\nu)\},$ (3.5)
+where $\tau_{n}$ is a tuning parameter that we set in accordance with the literature, [^12] Specifically, $\tau_{n}=\sqrt{\log(\underline{n})/\underline{n}}$, where $\underline{n}=\tfrac{2qn}{k(k+1)}$ is expected sample cell size for any but the universal choice problem; here, $k$ is the number of nondefault items in $X$ and $q$ is the number of “small” choice problems faced by each subject.
+Recall also that $H$ is the length of $\nu$, thus division by $H$ ensures that the above constraint is scaled by $\tau_{n}$ and not by the testing problem’s complexity.
+$\bm{1}$ is a vector of $1$’s, and $\hat{\pi}^{*}$ is a simple nonparametric (clustered, as explained earlier) bootstrap analog of $\hat{\pi}$.
+We go beyond a completely straightforward implementation of Kitamura and Stoye’s (2018) test because we do not weight questions equally.
+This possibility is anticipated by Kitamura and Stoye (2018), who only a diagonal weighting matrix $\Omega$, but has not, to our knowledge, been implemented before.
+We use it because, in our data, choice probabilities pertaining to the universal choice set $X$ will be estimated from a much larger sample cell than others.
+We take this into account by weighting estimated probabilities for different questions by the expected sample cell size for that question; in practice, that means to weight all small questions equally and to put weight $w\approx 9$ on choice frequencies corresponding to the grand set.
+[^13] More precisely, $w=\tfrac{k(k+1)}{2q}$, with $(k,q)$ as in the previous footnote.
 We do not weight questions by realized sample cell sizes, and we also do not estimate cell-specific variances by the binomial variance formula, in order to avoid data dependent weighting.
-In our application, these modifications would have minimal effect.
-We close by observing that in general, this test can be expensive to compute.
-The experimental design that we settled on, partly to ensure reasonable sample cell sizes, is small enough so that this is not the case.
+In our empirical application, these modifications would have minimal effect.
+In general, this test can be expensive to compute.
+The experimental design that we settled on, partly to ensure reasonable sample cell sizes, is small enough so that this concern does not arise.
 However, in preparation, we also implemented an adaptation of the computational improvements in Smeulders et al.
 (2021).
-For the benefit of future users, these details are laid out in Appendix B.
+These details are laid out in Supplementary Appendix B.
 4 An Application to Experimental Data
-We next describe an application of the above methods to an experimental data set.
 4.1 Experimental Design
-The aim of the experiment is to collect data of the type needed to implement tests for the Min and RUM bounds.
-This means we require repeated observations of choices from a grand set of alternatives and a number of subsets, with all sets containing a default option.
-Moreover, based on the findings of Chernev et al.
+Our testing strategy requires observing choices from a grand set of alternatives and a number of subsets.
+Based on the findings of Chernev et al.
 (2015), we want the choice problems to be non-trivial to increase the probability of finding choice overload.
-To this end, we ask subjects to make choices between amounts of experimental points expressed as sums, as in Caplin et al.
-(2011).
-Specifically, each non-default option is expressed as a sum of four numbers between zero and ten written in text.
+To this end, we ask subjects to make choices between amounts of experimental points expressed as sums, where each non-default option is expressed as sum of four numbers between 00 and $10$ written in text.
 The value of choosing an option in experimental points is the value of the sum, and one experimental point is worth $50$ cents.
 There are $12$ non-default options in the grand set $X$.
-Following Caplin et al.
-(2011), we generated these as follows: First, we drew the value of an alternative was drawn from an exponential distribution truncated at 10 points with $\lambda=0.25$.
-Next, individual terms of sums were chosen stochastically ensuring neither the first nor the maximal summand was correlated with the total value of the option.
-Meanwhile, a default option was available in each choice set.
-This option provided 7 points and was the only option expressed as a single number.
-The default option also always appeared at the top of the screen, was pre-selected (i.e.
-the subject had to actively click on another option if they wanted to choose it) and the simplest alternative (as measured by the number of summands in the expression).
-Thus it fulfilled many criteria that have been proposed to define the “default” in choice overload situations (see for example Iyengar and Kamenica (2010)).
-Of the non-default options, 9 yielded a number of points strictly lower than the default, 1 yielded the same number of points as the default, and 2 yielded strictly more points.
-Figure 1 shows an example choice screen with the default and one other option.
-Figure 2 shows an example choice screen for the grand set.
-Figure 1: A choice problem comparing one lottery to the default.
-Figure 2: A choice problem comparing the grand set to the default.
-The collection of choice sets used in the experiment consisted of the grand set (12 alternatives plus the default), and all 1 and 2 alternative subsets (along with the default), for a total of 78 smaller choice sets.
-Based on the prior literature, we initially believed that 3 item sets are unlikely to trigger choice overload, while 13 item sets are likely to do so, if such an effect is present.
+We generate these by first drawing the value of an alternative from an exponential distribution truncated at 10 points with $\lambda=0.25$, then randomly selecting individual terms of the sum so that neither the first nor the maximal summand was correlated with the total value of the option.
+(All of this follows Caplin et al.
+(2011).)
+Each choice set also contained a default option providing $7$ experimental points, expressed as a single number.
+This option furthermore appeared at the top of the screen and was pre-selected, so it was an obvious default choice.
+Of the non-default options, $9$ yielded a number of points strictly lower than the default, $1$ yielded the same number of points, and $2$ yielded strictly more points.
+Figure LABEL:fig:1(a) shows an example choice screen with the default and one other option.
+Figure LABEL:fig:1(b) shows an example choice screen for the grand set.
+The collection of choice sets consisted of the grand set and all subsets containing 1 or 2 alternatives along with the default, for a total of 78 smaller choice sets.
+Based on the prior literature, we initially believed that 3-item sets are unlikely to trigger choice overload, while 13-item sets are likely to do so if such an effect is present to begin with.
 Each subject was presented with 9 randomly selected small sets and the grand set, with the order of choice questions and the order of non-default options in each choice question randomized.
@@
 Subjects in addition received a \$1 participation fee.
-A complete list of the choice alternatives and choice sets can be seen in Appendix A.2.181818The experiment was approved by the Institutional Review Boards of Columbia and Cornell Universities.
-Note that we have selected a choice environment that should generate a good deal of heterogeneity in item quality, but in which there should be little preference heterogeneity.
-Based on the discussion of Section 3.1 we would anticipate the Min bound to be effective in identifying choice overload, whle the RUM bound may not necessarily provide further advantages.
+A complete list of choice alternatives and choice sets is in Supplementary Appendix A.2.141414The experiment was approved by the Institutional Review Boards of Columbia and Cornell Universities.
+we note that this is a situation in which we would anticipate item heterogeneity to outweigh preference heterogeneity – assuming people like more money to less, there should not be any preference heterogeneity.
+This means that, as per our discussion in Section 4.2, we would expect the Min bound to perform well relative to the RUM bound.
 Subject Recruitment.
-The experiment was run on Amazon’s Mechanical Turk (MT) platform, a digital marketplace for work.
-This platform was chosen to make it easy to get data from a large number of subjects, each of whom needed to answer only a limited number of questions.
-“Requesters” post Human Intelligence Tasks, or “HITs,” which are usually simple, repetitive jobs that typically pay small sums for each completed task.
+The experiment was run on Amazon’s Mechanical Turk (MT) platform.
+This platform was chosen to easily collect data from a large number of subjects, each answering a small number of questions.
+“Requesters” post Human Intelligence Tasks (“HITs”) – usually simple jobs that pay small sums for each completed task.
 Workers on MT view descriptions of the HITs, decide which to accept, and complete those HITs over the internet.
-In our case, subjects who accepted the HIT followed a link to an external webpage, where they completed the experiment.
-Upon completion they were given a randomly generated code, which was used to pay them the appropriate amount given their choices in the experiment.
-Instead of posting HITs and recruiting subjects directly on MTurk, we used an online platform called CloudResearch.
-CloudResearch posts our experiment as a HIT on MTurk and recruits the subjects.
-All subjects recruited through CloudResearch’s platform are first screened through CloudResearch’s Sentry data validation system.
-Sentry is a short (1 minute) pre-experiment survey meant to ensure subjects are attentive, engaged, and ready to participate in the experiment.
-People who fail Sentry are routed away from the experiment.
-These pre-study validation surveys have been shown to be effective in increasing data quality (Chandler et al., 2019; Litman and Robinson, 2020).
-We recruited 2000 subjects for the experiment.
+In order to improve subject attentiveness and reduce the probability of responses from bots, we prescreened subjects using the platform CloudResearch.
+This has been shown to be effective in increasing data quality (Chandler et al., 2019; Litman and Robinson, 2020).
+We recruited 2000 subjects in April and May of 2022.
 In addition to CloudResearch’s screening, we restricted recruitment to MTurk workers who had completed over 1000 HITs and had an approval rating of over 97%.
-Following the instructions subjects had to answer two quiz questions to ensure the understood the instruction.
-1833 subjects passed the quiz and completed the experiment, but one of these subjects had a browser related error which made their data unusable, leaving us with 1832 subjects’ data.
-The resulting data provides 1832 choices from the grand set.
-We have at least 185 observations of choices from each of the small sets, with variation in sample size due to the random selection of choice questions.
-4.2 Data Overview
-Figure 3 presents a histogram showing the distribution of choice sets by the probability with which the default option was chosen from sets other than $X$.
-The distribution is bimodal.
-The choice sets with low probability of default choice are precisely the ones in which the best option yields more points than the default.
-The probability of default choice in $X$ is indicated by the dotted vertical line.
-It falls somewhat above the default choice probability of the lower group, but below that of the higher group.
-Appendix A.2 lists the default choice probabilities for each choice set.
-Figure 3: Histogram showing distribution of choice sets by probability that default option was chosen. Dotted line shows the fraction of default choice in the grand set $X$.
+1833 subjects passed a comprehension quiz [^15] Screenshots of the quiz and instructions can be found in Supplemental Appendix A.1.
+and completed the experiment, but one of these subjects had a browser-related error which made their data unusable, leaving us with 1832 subjects’ data.
+In addition to 1832 choices from the grand set, we have at least 185 observations of choices from each of the small sets, with variation in sample size due to the random selection of choice questions.
+Supplementary Appendix A.2 lists default choice frequencies for each choice set.
 4.2 Analysis
-We next ask whether the data we collected are suggestive of choice overload by conducting tests in roughly increasing order of presumed power.
-First, we establish that existing approaches would be unlikely to identify choice overload in our data set.
-The empirical frequency of default choice in the grand set, $\hat{p}_{d}(X)$, was $22.3\%$.
-The analogous empirical frequency across all subsets was $70.7\%$ and therefore considerably higher.
-A simple comparison of means would not reveal choice overload, but also masks extreme heterogeneity across choice items.
-A random selection of a single smaller set would also be unlikely to reveal choice overload.
-$23$ of $78$ small choice sets had default choice frequency below $22.3\%$, only $15$ of which significantly so at the 5% level (using Fisher’s exact test on disjoint samples but without controlling for multiple comparisons).
-Thus, if a researcher were to randomly select a small choice set to compare to the grand set, they would only find evidence for choice overload 19% of the time.
-In a next step, we use the finite sample test to ask whether any of those $23$ frequencies is significantly below the analogous frequency in the grand choice set, taking into account sampling uncertainty.
-Strikingly, the answer is yes: The p-value against the null hypothesis that the grand choice default frequency is lowest is .00003.00003.00003.00003, and $5$ choice items are significantly below it at the $5\%$-level, where the p-values and significance levels just reported reflect Bonferroni adjustment across $78$ tests.
-[^19] The choice sets alluded to are numbers $(9,11,20,32,74)$ in order of appearance in Table 2.
+We test for choice overload in roughly increasing order of presumed test power.
+[^16] To showcase applicability of our methods to the data structure that we assumed, we ignore some features of our data at hand, namely that we observed precise choices and that we have several choices per subject, although we do consider the effect of the latter on sampling uncertainty.
+First, classic tests from the literature would be unlikely to find evidence of choice overload.
+The frequency of default choice in the grand set was $22\%$; the analogous frequency across all subsets was $71\%$.
+A simple comparison of means would, therefore, not reveal choice overload, and a random selection of a single smaller set would also be unlikely to do so – only 15 of 78 (19%) small sets have default choice significantly below the grand set’s.
+Next, we use the finite sample test to ask whether any single default frequency is significantly below the one in the grand choice set, taking into account sampling uncertainty and multiple testing.
+Strikingly, the answer is yes: After Bonferroni adjustment, the p-value against the null hypothesis that the grand choice set default probability is lowest is .00003.00003.00003.00003, and $5$ choice items are significantly lower at the $5\%$-level.
 Similarly, the asymptotic test of the Min bound yields a p-value that we could not distinguish from 00 in $B=10000$ Monte Carlo simulations.
-Hence we find strong evidence of choice overload using this approach.
-We next apply the test explained in Section 3.2.3 to see if our data is consistent with RUM.
-At a Monte Carlo replication size of $B=10000$, the p-value against this null hypothesis is equals 00 as well.
-This extremely strong rejection comes with a caveat: The p-value against all but the grand choice set equals .005.005.005.005.
+Hence, we find strong evidence of choice overload using this approach.
+We next test consistency with RUM.
+At a Monte Carlo replication size of $B=10000$, the p-value equals 00 as well.
+This strong rejection comes with a caveat: The p-value against all but the grand choice set equals .005.005.005.005.
 Therefore, “the data are consistent with RUM except that default choice from $X$ is too frequent” is not an appropriate description of our findings.
-However, this leaves open the possibility that choice overload had an effect already in some of the smaller choice sets.
+However, this leaves open the possibility that, contrary to our own prior, choice overload had an effect in some of the smaller choice sets.
 […]
@@
 $\max\bm{1}^{\prime}\nu\text{ s.t. }\bm{A}\nu\leq\hat{\pi}$
-and equals $1$ if, and only if, sample proportions are rationalizable in terms of rational matrix $\bm{A}$.
-This fraction is .866.866.866.866 for the RUM, i.e.
-model (i) above, increases to .877.877.877.877 for model (ii), and equals .915.915.915.915 for model (iii).
+and equals $1$ if, and only if, sample frequencies are rationalizable in terms of the behavior encoded in $\bm{A}$.
+This fraction is .866.866.866.866 for the RUM, increases to .877.877.877.877 for model (ii), and equals .915.915.915.915 for model (iii).
 Third, one may worry that model (iii) is just not very restrictive, while maybe models (i) and (ii) are.
-This cannot be literally true because our test statistics are positive in all three tests, hence empirical choice frequencies are not rationalizable under any model.
-But it could certainly be “morally true,” notably if all possible data sets are “close” to the model.
-In order to shed some light on this, we conducted some analyses inspired by Bronars (1987), Selten (1991), and Beatty and Crawford (2011).
-Specifically, for models (i)-(iii) we calculated the expected mean square error (MSE) when matching data that were generated from the uniform distribution on $[0,1]^{79}$ (i.e.
-choice probabilities drawn uniformly at random).
+This cannot be literally true because our test statistics are positive in all three tests, hence empirical choice frequencies do not conform to any model.
+But it could be “morally true,” notably if all possible data sets are close to the model.
+We investigated this through analyses inspired by Bronars (1987), Selten (1991), and Beatty and Crawford (2011).
+Essentially, these approaches propose comparing the test statistic observed in the data to that that observed in randomly generated pseudo data that matches the true data in some regards (for example, the choice sets from which choices are observed).
+Specifically, for models (i)-(iii) we calculated the expected mean square error (MSE) when matching data that were generated from the uniform distribution on $[0,1]^{79}$.
 The resulting values are 0.25 for model (i), 0.23 for (ii) and 0.21 for (iii).
-These numbers are all close to each other, and far above the MSE when these models are applied to the data (which range from 0.03 to 0.01), showing that even our most permissive model places significant restrictions on the data.
-Finally, the data can be used to illustrate the difference between the Min bound, $p_{d}^{min}(X)$, and the RUM-based bound, $p_{d}^{RUM}(X)$.
-The empirical choice frequencies do not imply a well-defined $p_{d}^{RUM}(X)$ because they are not stochastically rationalizable even if choice set $X$ is ignored.
-However, one can easily compute the vector $\hat{\eta}$ of choice probabilities that are closest to the empirical ones while being rationalizable; in fact, this computation is a by-product of the statistical tests.
+These numbers are all close to each other and far above the MSE when these models are applied to the data (which range from 0.03 to 0.01), showing that even our most permissive model places significant restrictions on the data.
+Finally, the data can be used to illustrate the difference between the Min and RUM bounds.
+While empirical choice frequencies do not imply a well-defined $p_{d}^{RUM}(X)$, one can easily compute the vector $\hat{\eta}$ of choice probabilities that are closest to the empirical ones while being rationalizable; in fact, this computation is a by-product of the statistical tests.
 This allows us to compute a feasible analog of $p_{d}^{RUM}(X)$.
@@
 To compare apples to apples, we report that the same $\hat{\eta}$ implies a Min bound of $p_{d}^{min}(X)=11.4\%$.
-Using the full implications of RUM is potentially much more informative than just testing monotonicity.
+Once again, using the full implications of RUM is potentially much more informative than just testing monotonicity.
 5 Conclusion
-In this paper, we have argued that existing tests for choice overload are underpowered, and that this may explain the confusing and contradictory picture that emerges from the current literature.
-We proposed that, by collecting more data and making use of theory, one can design better tests.
-We have demonstrated that, in a novel data set, our approach finds choice overload.
-Indeed, our new data speak so loudly that with hindsight, our methodological innovations would not have been necessary to detect choice overload in our own data set.
-We believe that the innovations are of interest nonetheless, and we also note that such things are bound to occur if –as we did– one genuinely designs the empirical strategy before going into the field.
-One unexpected finding from our experiment is that choice appears to violate the RUM even in smaller choice sets.
-This is perhaps surprising given the relative simplicity of the choice items.
-Understanding the cause of these violations strikes us as an interesting avenue for future research.
-For our purposes, it suggests that RUM may not be an appropriate benchmark against which to look for choice overload, at least in this experimental task.
+In this paper, we argued that existing tests for choice overload are not very sensitive, potentially explaining the ambiguous picture that emerges from the current literature.
+We proposed that, by collecting more data and fully using restrictions from economic theory, one can design better tests.
+We find choice overload in a novel data set, while standard approaches would only have had a 19% chance of doing so.
+Indeed, our new data speak so loudly that with hindsight, our econometric innovations would not have been necessary to detect choice overload.
+We believe that the innovations are of interest nonetheless, and we also note that such things are bound to occur if one genuinely designs the empirical strategy before collecting data (as we did).
+We hope that our work will have three consequences.
+First, by providing a higher powered test of choice overload, it should clear up the question of whether this is indeed a real phenomenon.
+Second, given that (we suspect) it will show choice overload to be more prevalent than previously thought, we hope it will spur further theoretical and policy work designed to understand its causes and mitigate its effects.
+Finally, by providing a better tool for measuring when choice overload does occur, we hope it will facilitate the above work by providing a better empirical basis on which to theorize.
 Appendix A Supplementary Appendix: Experimental Details
-Here we present the experiment instructions, quiz, and details about the choice objects.
 A.1 Instructions and Quiz
-Subjects were first shown this instructions screen before proceeding to the quiz (below). Subjects were given two attempts at the quiz.
-[Uncaptioned image] [Uncaptioned image]
+Subjects were first shown an instructions screen before proceeding to the quiz; see Figure 3. Subjects were given two attempts.
+Figure 3: Instructions.
+Figure 4: Quiz screen 1.
+Figure 5: Quiz screen 2.
+Figure 6: Quiz screen 3.
 A.2 Choice Alternatives with Summary Data
-Table 1 gives full list of the choice objects with their values in experimental points.
-Table 1: List of Options with Values
+Table 1 gives a full list of the choice objects with their values in experimental points.
+Table 2 shows how often each choice set was shown to subjects and how often the default was chosen from it.
+Choice sets are given by the alternatives they contain: e.g.
+$[0,1,2]$ represents the choice set with the default (0) and options with IDs 1 and 2 from Table 1.
+Figure 7 shows a histogram of the default choice probabilities of all choice sets.
 ID Option Value
@@
 12 six minus ten plus seven minus two 1
-Table 2 shows how often each choice set was shown to subjects and how often the default was chosen from it.
-Choice sets are given by the alternatives they contain: $[0,1,2]$ represents the choice set with the default (0) and options with IDs 1 and 2 from Table 1.
+Table 1: List of Options with Values
 Choice Set # Choices # Default Choice Set # Choices # Default
@@
 Table 2: Choice Set and Default Choice Frequencies
-Appendix B Implementing recent computational innovations
-We implemented the computational procedure in Smeulders et al.
+Figure 7: Histogram showing distribution of choice sets by probability that default option was chosen. Dotted line shows the fraction of default choice in the grand set $X$.
+Appendix B Supplementary Appendix:
+Implementing Recent Computational Innovations
+In the process of writing this paper, we implemented the computational procedure in Smeulders et al.
 (2021).
@@
 The implementation and some not entirely obvious modifications are described next.
-The procedure is motivated by the fact that computation of the matrix $\bm{A}$ and also computation (3.4) is hard, and yet the latter needs to be repeated many times.
+This material is not in the main text because we did not end up using the implementation in the empirical work.
+Smeulders et al.’s (2021) procedure is motivated by the fact that computation of the matrix $\bm{A}$ and also computation (3.4) is hard, and yet the latter needs to be repeated many times.
 It exploits that, because $\bm{BA}$ has many more columns than rows, there always exists a sparse (in the loose sense of having relatively few nonzero entries) $\arg\max$ to problem (3.4).